Jefferson Lab HALL A Logbook for June 10 - All EXCEPT Automatic Entries

  • Main INDEX

    Wednesday

  • 329239 06/30/10 23:29 J. Zhang Swing shift summary
  • 329238 06/30/10 23:28 J. Zhang Control access (Late entry)
  • 329227 06/30/10 22:06 J. Zhang Beam line test summary
  • 329226 06/30/10 22:05 J. Zhang Beam line test summary
  • 329225 06/30/10 21:53 J. Zhang Beam line test finished. Septum off
  • 329224 06/30/10 21:36 stepanyan target move
  • 329222 06/30/10 20:47 stepanyan target move
  • 329221 06/30/10 19:18 Stepanyan target move
  • 329220 06/30/10 18:15 J. Zhang Beam line test will be started soon
  • 329219 06/30/10 18:11 J. Zhang Ready for beam permit
  • 329218 06/30/10 17:25 rwright Magnet Interface to FSD Readiness
  • 329216 06/30/10 17:19 J. Zhang Hall closed
  • 329214 06/30/10 17:10 R. Michaels TDC problem, roc1 slot 11
  • 329213 06/30/10 17:07 R. Michaels trig / daq status update
  • 329212 06/30/10 16:49 ZAFAR AHMED First production run of PREX after ADCs gain changed
  • 329211 06/30/10 16:07 hsmith Right Dipole Polarity change
  • 329210 06/30/10 16:04 hsmith Quench circuit Right Dipole
  • 329208 06/30/10 16:02 LeRose Shift Summary
  • 329207 06/30/10 14:54 riordan Foils and Ta y vs phi
  • 329199 06/30/10 13:46 techs check list
  • 329198 06/30/10 12:00 R. Michaels trigger / daq status
  • 329197 06/30/10 11:39 ZAFAR AHMED Log out and log in adaql7
  • 329172 06/30/10 06:14 I Pomerantz The mouse in the backroom got caught in the glue-trap, was released unharmed near the south fence.
  • Tuesday

  • 329171 06/29/10 23:21 Markowitz Swing shift summary
  • 329162 06/29/10 19:42 ejensen Pion rejection efficiency for high rate
  • 329161 06/29/10 19:33 R. Michaels R-HRS laptop status
  • 329160 06/29/10 19:20 Jack Right VDCs gas flow
  • 329159 06/29/10 17:31 ellie Day Shift Summary
  • 329158 06/29/10 17:27 R. Michaels Several BCMs unplugged and related checks
  • 329157 06/29/10 17:11 riordan Present APEX Run Plan
  • 329155 06/29/10 16:13 V. Sulkosky BPM Calibration Completed
  • 329153 06/29/10 15:56 tecs ckecklist
  • 329152 06/29/10 14:35 schuster target to optics
  • 329151 06/29/10 14:32 ellie Small, erratic Q3 fluctuations
  • 329141 06/29/10 14:02 ellie Hall A Tools Screen Grab
  • 329138 06/29/10 13:29 Ole Hansen VDc efficiencies run 1336
  • 329137 06/29/10 13:22 Ole Hansen VDC 1uA vs 32uA high-rate running observations
  • 329132 06/29/10 11:51 ellie Slight rise in target temp and vacuum
  • 329128 06/29/10 11:13 ellie Changing prescales to 1 for cosmics run
  • 329127 06/29/10 09:19 ellie Hall is now in Restricted Access
  • 329124 06/29/10 07:59 Jin Huang Shift Summary : June 29 Owl
  • 329123 06/29/10 07:46 kelleher Septum is OFF
  • 329121 06/29/10 07:41 Jin Huang All tests done. Hall -> Restricted Access
  • 329118 06/29/10 07:31 Jin Huang Q3 current oscillating @ 500MeV/c
  • 329115 06/29/10 07:21 Jin Huang LHRS -> 0.700GeV/c -> 0.500GeV/c
  • 329108 06/29/10 06:50 Jin Huang HRS->0.900GeV/c
  • 329099 06/29/10 06:29 Jin Huang LHRS -> 1.000GeV/c
  • 329086 06/29/10 06:08 Jin Huang Optics Run Plot
  • 329081 06/29/10 05:52 Jin Huang Target -> Optics and Hall A tools
  • 329079 06/29/10 05:37 Jin Huang BPM Calibration: A=B=(-3,+3), Raster OFF
  • 329076 06/29/10 05:30 Jin Huang BPM Calibration: A=B=(-3,-3), Raster OFF
  • 329074 06/29/10 05:21 Jin Huang BPM Calibration 2: A=B=(+3,-3), Raster OFF
  • 329068 06/29/10 05:01 Jin Huang BPM Calibration 3: A=B=(+3,+3), Raster OFF
  • 329065 06/29/10 04:47 Jin Huang BPM Calibration 2: A=B=(0,0), Raster OFF
  • 329062 06/29/10 04:38 Jin Huang BPM Calibration 1: A=B=(0,0), Raster = 10x5MCC
  • 329060 06/29/10 04:31 eip targte change
  • 329057 06/29/10 04:24 Jin Huang Parameters for 32uA run 1336
  • 329054 06/29/10 04:09 Jin Huang Parameters for 24uA run 1335
  • 329051 06/29/10 04:03 Jin Huang VDC HV current VS beam current
  • 329050 06/29/10 03:52 S.Abrahamyan GC gap in acceptance
  • 329045 06/29/10 03:41 kelleher target: moved to Pb target
  • 329043 06/29/10 03:34 Jin Huang Moved to Ta target for MCC radiation calibration
  • 329041 06/29/10 03:29 I Pomerantz TOF s0 <-> cerenkov
  • 329040 06/29/10 03:29 eip target change
  • 329038 06/29/10 02:50 kelleher target: moved to empty target
  • 329031 06/29/10 02:34 Jin Huang Spot++, 10x5 raster
  • 329028 06/29/10 02:31 Ole Hansen VDC x and xy dists for evt5
  • 329027 06/29/10 02:30 Jin Huang Spot++, 8x5 raster
  • 329022 06/29/10 02:06 Jin Huang Spot++, 5x5 MCC raster on lead, 1uA
  • 329018 06/29/10 01:58 S.Abrahamyan S2m ADC mean value
  • 329017 06/29/10 01:54 kelleher Target Moved: Lead #1
  • 329013 06/29/10 01:42 I pomerantz Cerenkov TDC pairs
  • 329010 06/29/10 01:30 mgraham accidental rate for VDC
  • 329008 06/29/10 01:27 S.Abrahamyan Gas Cherenkov mean value
  • 329005 06/29/10 01:19 N Toro relative trigger rates ok
  • 328999 06/29/10 01:07 kelleher target move: Ta
  • 328998 06/29/10 01:06 Jin Huang Q2/3 Cycle done
  • 328997 06/29/10 00:45 Jin Huang Tune Beam in Hall
  • 328996 06/29/10 00:45 kelleher target motion: empty target
  • 328995 06/29/10 00:41 Jin Huang Cycling Q2/3 on LHRS. They were not cycled since polarity flip!
  • 328993 06/29/10 00:27 mgraham multiple peaks in tdc
  • 328991 06/29/10 00:12 N Toro thin carbon scaler rates
  • 328987 06/29/10 00:10 eip tragte change
  • 328985 06/29/10 00:06 Ole Hansen VDC time-to-distance
  • 328984 06/29/10 00:05 J. Zhang Swing Shift Summary
  • Monday

  • 328981 06/28/10 23:57 Stepanyan target move
  • 328980 06/28/10 23:55 N Toro very different behavior for Pb, Ta trigger rates
  • 328978 06/28/10 23:53 J. Zhang Ta target, 0.5 micro A
  • 328976 06/28/10 23:49 N Toro scaler rates 0.5uA on Ta
  • 328975 06/28/10 23:48 N Toro scaler rates 0.5uA on Ta target
  • 328974 06/28/10 23:47 Stepanyan target move
  • 328971 06/28/10 23:35 Stepanyan target move
  • 328968 06/28/10 23:29 J. Zhang rate with BeO target
  • 328967 06/28/10 23:26 S.Abrahamyan Gas Cherenkov mean value
  • 328966 06/28/10 23:16 S.Abrahamyan Gas Cherenkov
  • 328965 06/28/10 23:13 S.Abrahamyan Gas Cherenkov
  • 328962 06/28/10 23:05 Stepanyan target is BeO
  • 328961 06/28/10 23:03 J. Zhang Change to BeO target
  • 328960 06/28/10 23:02 N Toro spot++ 8x4 mcc
  • 328959 06/28/10 22:55 J. Zhang Continue the beam recovery procedure
  • 328956 06/28/10 22:50 Jack (Techs) Septum magnet power supply breaker tripped
  • 328954 06/28/10 22:41 J. Zhang Go back to beam recovery
  • 328944 06/28/10 22:14 J. Zhang Escorted acess
  • 328943 06/28/10 22:11 J. Zhang Septum IOC rebooted
  • 328941 06/28/10 21:33 N Toro spot++ run 1306
  • 328938 06/28/10 21:26 J. Zhang Septum read back currents are wrong
  • 328937 06/28/10 21:20 N Toro 1305 spot++
  • 328932 06/28/10 21:07 N Toro spot++ run 1303
  • 328931 06/28/10 21:04 adaq Spot++ test
  • 328928 06/28/10 20:17 J. Zhang Septum on
  • 328927 06/28/10 20:10 J. Zhang Move target to empty
  • 328926 06/28/10 20:10 J. Zhang Hall closed, Beam recovery started
  • 328925 06/28/10 20:04 riordan Late Entry: Hall in Controlled Access
  • 328923 06/28/10 17:44 Folts checklist/tech
  • 328922 06/28/10 15:46 Fansler/Ewin Polarity change
  • 328921 06/28/10 15:18 R. Michaels new DAQ configs: TWOARM, etc
  • 328910 06/28/10 11:50 R. Michaels spot++ fixed (gate polarity was wrong)
  • 328901 06/28/10 11:25 R. Michaels cleanup of 1024 Hz pulse
  • 328900 06/28/10 11:00 R. Michaels ROC4, slot 24 gate problem
  • 328899 06/28/10 09:49 S. Abrahamya correction
  • 328898 06/28/10 09:47 S. Abrahamya Pion rejection.
  • 328897 06/28/10 09:29 S. Abrahamya Gas Cherenkov
  • 328896 06/28/10 09:16 S. ABrahamya Pion rejector
  • 328895 06/28/10 09:09 V. Sulkosky 1024 Hz Clock Signal BAD (?)
  • 328894 06/28/10 08:17 Markowitz VDC current readout not working
  • 328892 06/28/10 07:55 Markowitz Septum off for access
  • 328891 06/28/10 07:54 Markowitz Owl shift summary: June 28, 2010
  • 328890 06/28/10 07:53 adaq tools
  • 328888 06/28/10 07:45 S. Abrahamya GC HV
  • 328885 06/28/10 07:36 S. Abrahamya GC HV
  • 328881 06/28/10 07:26 S.Abrahamyan GC HV
  • 328880 06/28/10 07:24 kelleher Target: Moved to Pb/Diamond
  • 328879 06/28/10 07:17 kelleher Target: Moved to Thick C
  • 328876 06/28/10 07:08 kelleher Target: Moved to Pb Diamond #1
  • 328875 06/28/10 07:07 S. Abrahamya GC HV
  • 328872 06/28/10 06:50 A. Camsonne Spot check request on Pb target
  • 328871 06/28/10 06:40 kelleher Target: moved to C
  • 328868 06/28/10 06:35 kelleher Target to Pb Diamond #1
  • 328867 06/28/10 06:31 S.Abrahamyan Gas cherenkov HV
  • 328866 06/28/10 06:30 Markowitz rates with and without beam
  • 328863 06/28/10 06:22 Markowitz Moving to thick carbon target
  • 328860 06/28/10 05:51 Markowitz Septum back on
  • 328859 06/28/10 05:06 S. Abrahamya Pion rejector
  • 328858 06/28/10 04:53 Markowitz More Q1 oscillations
  • 328857 06/28/10 04:45 Markowitz Septum off for access
  • 328852 06/28/10 04:21 markowitz move to Pb-D1
  • 328847 06/28/10 03:59 Markowitz move to Ta
  • 328844 06/28/10 03:49 Markowitz Septum back on
  • 328836 06/28/10 03:15 adaq Return of the Q1 oscillation
  • 328828 06/28/10 02:56 Markowitz Call for controlled access, turn off septum
  • 328827 06/28/10 02:51 Markowitz Call for controlled access, turn off septum
  • 328826 06/28/10 02:47 Markowitz Move to thick carbon
  • 328816 06/28/10 02:20 Markowitz putting in Tantalum target
  • 328815 06/28/10 02:12 Markowitz First beam in hall
  • 328814 06/28/10 01:19 V. Sulkosky Restored Clock to L-HRS
  • 328810 06/28/10 00:22 mgraham additional PiR channels
  • 328806 06/28/10 00:06 Jin Huang Shift Summary : June 27 Swing
  • Sunday

  • 328804 06/27/10 22:10 mgraham gc tdc noise as expected
  • 328803 06/27/10 22:08 Jin Huang Control Access
  • 328802 06/27/10 21:00 Jin Huang Q1 current was unstable, solved
  • 328797 06/27/10 20:50 A. Camsonne Gas cerenkov rates with no beam
  • 328794 06/27/10 20:18 Jin Huang Left Q1 Fault, Recovered
  • 328791 06/27/10 19:52 Jin Huang Beam Tuning Status
  • 328788 06/27/10 19:30 mgraham gc HV changed back to default
  • 328784 06/27/10 19:22 mgraham temporary gc hv change
  • 328782 06/27/10 19:13 A. Camsonne Requested to go back to beam permit
  • 328779 06/27/10 18:47 mgraham cerenkov tdc noisy
  • 328774 06/27/10 18:28 A. Camsonne MCC still tuning for Hall A - access possible
  • 328773 06/27/10 18:25 V. Sulkosky Disabled Pedestal Suppression for L-HRS
  • 328770 06/27/10 18:05 Yi Qiang Gas Cherenkov ADC finally shows up
  • 328767 06/27/10 17:56 mgraham cerenkov adc histograms
  • 328766 06/27/10 17:46 Jin Huang Beam update
  • 328764 06/27/10 17:42 hsmith Left MNR repalcemet
  • 328763 06/27/10 17:31 hsmith Left dipole polarity switch
  • 328762 06/27/10 17:27 Jin Huang LHRS -> 1.131 GeV
  • 328761 06/27/10 17:26 Jack (Techs) Status of NMR readouts
  • 328758 06/27/10 17:12 Iqbal Target changed
  • 328757 06/27/10 16:51 R. Lindgren shift summary
  • 328755 06/27/10 16:33 mgraham similar behavior for corrected cerenkov amplitudes
  • 328754 06/27/10 16:06 Camsonne/Iqb Left Q2 & Q3 cycled
  • 328753 06/27/10 15:40 hsmith Left Dipole MNR
  • 328752 06/27/10 15:34 hsmith Left Diplole and Septem polarity switch
  • 328751 06/27/10 14:11 Richard Lind Hall A Tools
  • 328750 06/27/10 13:38 Jin Huang Optics Check : Sieve OUT runs
  • 328749 06/27/10 13:18 A. Camsonne Gas cerenkov HV off
  • 328748 06/27/10 12:48 A. Camsonne Q1,Q2 and Q3 polarity changed
  • 328747 06/27/10 12:33 A. Camsonne Q2 polarity change
  • 328746 06/27/10 12:25 Lindgren Hall A going to control access
  • 328745 06/27/10 12:24 A. Camsonne Ramping down magnets for polarity change / asked for controlled access
  • 328744 06/27/10 12:23 A. Camsonne HRS setting at 0.581 GeV/c
  • 328739 06/27/10 11:47 Lindgren -10% scan completed
  • 328735 06/27/10 11:35 A. Camsonne Short term plan of APEX test run
  • 328734 06/27/10 11:25 R. Michaels re: negative deadtime
  • 328732 06/27/10 11:22 R. Lindgren Starting 10% scan
  • 328728 06/27/10 11:05 LIndgren NMR Lock unstable
  • 328725 06/27/10 10:40 A. Camsonne Experiment status
  • 328718 06/27/10 10:28 ejensen Spot ++
  • 328715 06/27/10 10:24 Al Odian Move from BeO to Optics
  • 328713 06/27/10 10:17 ejensen Spot ++
  • 328704 06/27/10 10:00 A. Camsonne Moving to Emptry target
  • 328699 06/27/10 08:57 R. Lindgren Day shift starting waiting for beam
  • 328698 06/27/10 08:12 benesch operational limits
  • 328696 06/27/10 07:56 Markowitz Owl shift summary: June 27, 2010
  • 328695 06/27/10 07:46 S. Abrahamya Gas Cherenkov time
  • 328694 06/27/10 07:22 S.Abrahamyan changes due to gate change
  • 328693 06/27/10 07:21 ipom try again with line breaks
  • 328692 06/27/10 07:18 ipom s2m amplitudes allignment
  • 328691 06/27/10 06:39 R. Barbieri moving to optics target
  • 328688 06/27/10 06:29 Markowitz Left HRS set to 1.1383
  • 328687 06/27/10 06:20 R. Barbieri temp/vacuum chart
  • 328686 06/27/10 06:19 R. Barbieri move to tantalum target
  • 328685 06/27/10 06:17 Markowitz cycling quads
  • 328678 06/27/10 05:51 Markowitz Turn septum back on
  • 328677 06/27/10 05:50 Markowitz Request access
  • 328676 06/27/10 05:49 bogdanw Detector work
  • 328675 06/27/10 05:48 ipom HV adjusment for s2m
  • 328673 06/27/10 05:30 N Toro datamon reports negative deadtime
  • 328669 06/27/10 05:03 R. Barbieri temp/vacuum chart
  • 328668 06/27/10 04:57 S. Abrahamya Pion rejector status
  • 328661 06/27/10 04:40 R. Barbieri switch to thick carbon target
  • 328660 06/27/10 04:40 ipom HV change for s2m
  • 328657 06/27/10 04:25 R. Barbieri Moving to optics target.
  • 328649 06/27/10 04:05 Markowitz Magnets for detector commissioning
  • 328647 06/27/10 04:04 R. Barbieri temp/vacuum chart
  • 328643 06/27/10 03:53 Markowitz Should have been 1.1383, not 1.383 GeV/c
  • 328642 06/27/10 03:48 Markowitz Change momentum to 1.383 GeV/c
  • 328637 06/27/10 03:34 R. Barbieri Switch to Tantalum target
  • 328634 06/27/10 03:22 ipom Gas Cerenkov HV change
  • 328633 06/27/10 03:18 ipom HV changes to s2m
  • 328632 06/27/10 03:15 Markowitz Delta P = 0%, magnets for run 1212
  • 328631 06/27/10 03:07 Markowitz Reset Q2
  • 328630 06/27/10 03:05 R. Barbieri temp/vacuum stripchart
  • 328629 06/27/10 02:50 N Toro epicsRunBegin flags changed
  • 328625 06/27/10 02:31 N Toro modified runend.tcl
  • 328620 06/27/10 02:03 R. Barbieri target vacuum/temp normal
  • 328606 06/27/10 01:06 Markowitz tools screen
  • 328605 06/27/10 01:01 Markowitz LRSH set to 800 MeV/c for detector work
  • 328604 06/27/10 00:53 A. Camsonne Left HRS Sieve out
  • 328603 06/27/10 00:53 Ole Hansen VDC TDC offset calibration
  • 328602 06/27/10 00:43 kelleher Pion Rejector HV
  • Saturday

  • 328600 06/26/10 23:56 Yi Qiang Swing shift summary: June 26, 2010
  • 328599 06/26/10 23:54 Yi Qiang Septum magnets off for controlled access
  • 328598 06/26/10 23:52 Yi Qiang Controlled access to remove sieve slits
  • 328593 06/26/10 23:40 S. Abrahamya PR HV
  • 328590 06/26/10 23:30 mgraham gas cerenkov #2 HV
  • 328588 06/26/10 23:27 kelleher run 1197 -- Pion Rejector HV change
  • 328586 06/26/10 23:23 Yi Qiang Summary of runs for optics calibration
  • 328585 06/26/10 23:21 A. Camsonne Multiple peak in S2m timing spectrum
  • 328580 06/26/10 23:12 A. Camsonne Added VDC threshold set value to EPICS run start
  • 328573 06/26/10 22:57 Yi Qiang Spot++ of run 1194 with raster on
  • 328570 06/26/10 22:53 Ole Hansen End of run comments incorrect
  • 328566 06/26/10 22:46 Yi Qiang Left HRS p0 = 800 MeV/c
  • 328565 06/26/10 22:39 Yawei Target moved to optics!
  • 328559 06/26/10 22:34
  • 328558 06/26/10 22:32 A. Camsonne Right arm HV GUI restarted
  • 328554 06/26/10 22:25 Yawei Target moved to Ta position!
  • 328550 06/26/10 22:18 Yi Qiang Spot++ for run 1188 with raster ON
  • 328543 06/26/10 22:06 Yi Qiang Typo of momentum, should be 1.1266 GeV/c
  • 328542 06/26/10 22:05 Yi Qiang Left HRS p0 = 1.115 GeV/c
  • 328536 06/26/10 21:51 Yawei Target moved to optics!
  • 328534 06/26/10 21:45 Yi Qiang Left HRS p0 = 1.266 GeV/c before run 1182
  • 328529 06/26/10 21:33 Jin Huang Optics plot, dp = -1% & -2%
  • 328525 06/26/10 21:25 Yawei Target moved to Ta position!
  • 328523 06/26/10 21:18 kelleher Pion Rejector HV, part 2
  • 328522 06/26/10 21:18 Yi Qiang Spot++ of run 1178 with raster on
  • 328518 06/26/10 21:13 Jin Huang New replay database dir: /adaqfs/halla/adaq/happexsp/prex/DB/20100626
  • 328516 06/26/10 21:11 kelleher Pion Rejector HV
  • 328512 06/26/10 21:05 Yi Qiang Left HRS p0 to 1.1383 GeV/c
  • 328505 06/26/10 20:55 A. Camsonne Lecroy HV screen back
  • 328504 06/26/10 20:54 Yawei Target moved to optics!
  • 328503 06/26/10 20:52 Yi Qiang Left HRS po to 1.15 GeV/c
  • 328499 06/26/10 20:40 Jin Huang More Optics Stuff
  • 328497 06/26/10 20:36 kelleher Run #1166 VDC wires
  • 328492 06/26/10 20:29 A. Camsonne iocha22 rebooted ok but screens for HV still blank
  • 328491 06/26/10 20:27 Yawei Target moved to Ta position!
  • 328487 06/26/10 20:14 A. Camsonne iocha22 down - doing reboot from portserver
  • 328482 06/26/10 19:50 Jin Huang Spot++, 8x8, Run 1163
  • 328481 06/26/10 19:49 Jin Huang spot++ 7x7MCC, Run 1162
  • 328480 06/26/10 19:44 Yi Qiang Start to take optics data from run 1165!
  • 328471 06/26/10 19:14 kelleher raster check
  • 328469 06/26/10 19:08 Yawei Target moved to optics
  • 328467 06/26/10 19:02 V. Sulkosky Raster Check (5 x 5)
  • 328465 06/26/10 18:58 Yi Qiang Raster 5x5 on BeO target
  • 328464 06/26/10 18:56 V. Sulkosky Spot++ Raster Off
  • 328459 06/26/10 17:28 Yawei Vacuum plot
  • 328458 06/26/10 17:14 V. Sulkosky 50 ns delay Added to Fastbus TDC Gates
  • 328457 06/26/10 17:12 Yawei Target moved to BeO
  • 328456 06/26/10 17:11 Yi Qiang Beam is in the Hall, harp failed
  • 328455 06/26/10 16:53 Yi Qiang Set left HRS p0 to 1.16 GeV
  • 328453 06/26/10 16:26 V. Sulkosky Re: Day Shift Summary
  • 328452 06/26/10 16:20 V. Sulkosky Day Shift Summary
  • 328449 06/26/10 16:16 Yawei Target moved to empty position
  • 328447 06/26/10 15:43 Ole Hansen Please log run lists in shift summaries
  • 328433 06/26/10 14:55 R. Michaels datamon and deadtime updates
  • 328431 06/26/10 14:38 N Toro target cooling; controlled access
  • 328429 06/26/10 14:31 R. Michaels tests of scripts, bizarre need to ssh or GUIs dont work
  • 328410 06/26/10 14:10 R. Michaels runstart / runend scripts updated
  • 328407 06/26/10 13:58 Bartlett/Nei Reconnected wiring for BCMs in L-HRS hut(done 06/25/10)
  • 328406 06/26/10 13:57 V. Sulkosky Re: BPM/raster to roc4, slot 24
  • 328402 06/26/10 13:52 V. Sulkosky Added L1A signal to Norm. Scaler
  • 328400 06/26/10 13:26 ejensen 3rd screenshot
  • 328399 06/26/10 13:22 ejensen New pion rejector HV settings
  • 328397 06/26/10 13:20 R. Michaels BPM/raster to roc4, slot 24
  • 328395 06/26/10 13:11 ejensen 3rd screenshot of PR HV settings
  • 328394 06/26/10 13:07 Ole Hansen VDC timing instability?
  • 328392 06/26/10 13:04 ejensen Pion Rejecter HV adjustment
  • 328390 06/26/10 12:55 ejensen Gas Cherenkov HV adjustment
  • 328388 06/26/10 12:42 Bartlett/Nei T8 trigger Restored
  • 328383 06/26/10 12:09 Jin Huang Q2/Q3 cycle done
  • 328382 06/26/10 11:55 V. Sulkosky LHRS Network Switch Cycled
  • 328381 06/26/10 11:50 Jin Huang Cycling Q2/3 while Settling down Dipole
  • 328378 06/26/10 11:12 V. Sulkosky Septum Off
  • 328374 06/26/10 10:04 Jin Huang Optics Check out
  • 328373 06/26/10 10:03 V. Sulkosky Target Moves
  • 328371 06/26/10 10:00 N Toro planned access at noon
  • 328369 06/26/10 09:53 N Toro prescale not being logged
  • 328365 06/26/10 09:25 V. Sulkosky Septum back on
  • 328364 06/26/10 09:18 Jin Huang LHRS Sieve In
  • 328363 06/26/10 09:16 V. Sulkosky P0 Max Set point: 1.16 GeV/c
  • 328361 06/26/10 09:15 V. Sulkosky Septum Turned Off
  • 328360 06/26/10 08:57 V. Sulkosky Access to insert sieves
  • 328357 06/26/10 08:44 V. Sulkosky Hall A Tools (delta = 0%)
  • 328354 06/26/10 08:26 R. Barbieri optics target
  • 328350 06/26/10 08:13 R. Michaels how to fix spot++
  • 328348 06/26/10 08:09 Markowitz Day shift plan
  • 328347 06/26/10 08:06 bogdanw HRS-L detector work
  • 328344 06/26/10 08:00 Markowitz Owl shift summary
  • 328343 06/26/10 07:53 aserg VDC with beam
  • 328340 06/26/10 07:50 Markowitz cycling quads
  • 328339 06/26/10 07:41 R. Barbieri moving to empty target; harp scan
  • 328327 06/26/10 07:04 R. Barbieri vacuum plot
  • 328312 06/26/10 05:45 Markowitz Hall A access
  • 328311 06/26/10 05:34 R. Barbieri Vaccum + beam chart
  • 328293 06/26/10 04:22 vmaxwell VDC v2 wire and time run # 1097
  • 328288 06/26/10 04:16 R. Barbieri High voltage tuning
  • 328285 06/26/10 04:11 vmaxwell VDC v1 wire and time run # 1097
  • 328284 06/26/10 04:10 R. Barbieri Printing problem
  • 328283 06/26/10 04:08 vmaxwell VDC u2 wire and time run # 1097
  • 328282 06/26/10 04:04 vmaxwell VDC u1 wire and time run # 1097
  • 328281 06/26/10 04:03 Ole Hansen VDC checkout
  • 328279 06/26/10 03:56 R. Barbieri move to BeO
  • 328278 06/26/10 03:55 R. Barbieri mistaken move to BeO - back to empty
  • 328274 06/26/10 03:39 R. Barbieri moving to BeO target.
  • 328273 06/26/10 03:22 Ole Hansen VDC TDC calibration HOWTO
  • 328272 06/26/10 03:17 Markowitz magnet screens
  • 328271 06/26/10 03:06 Markowitz Left HRS set to 1 GeV/c.
  • 328269 06/26/10 03:01 Markowitz MCC update
  • 328267 06/26/10 02:55 vmaxwell run # 1095 histograms
  • 328266 06/26/10 02:42 R. Barbieri Vacuum, beam plots
  • 328265 06/26/10 02:39 Markowitz MCC update
  • 328260 06/26/10 01:40 vmaxwell online GUI histograms run # 1093
  • 328259 06/26/10 01:38 vmaxwell online GUI histograms run # 1093
  • 328258 06/26/10 01:33 vmaxwell Online GUI histograms run # 1093
  • 328257 06/26/10 01:03 Markowitz MCC update
  • 328256 06/26/10 00:44 A. Camsonne S0 HV set to 900 V
  • Friday

  • 328242 06/25/10 23:14 N Toro moving target to empty
  • 328241 06/25/10 23:09 A. Camsonne Plan for the next coming shift
  • 328238 06/25/10 22:31 brads (Partial) magnet temperature test
  • 328237 06/25/10 22:19 A. Camsonne Vacuum looking ugly
  • 328229 06/25/10 20:27 A. Camsonne Target vacuum snapshot
  • 328225 06/25/10 20:16 A. Camsonne Target vacuum
  • 328213 06/25/10 19:09 mgraham ADC slots, gates,and groups
  • 328210 06/25/10 18:26 Jack Status of Right Detector Stack
  • 328208 06/25/10 17:43 LeRose Shift Summary (Swing)
  • 328204 06/25/10 17:16 A. Camsonne Sweep scheduled for 8 PM
  • 328202 06/25/10 17:09 A. Camsonne Target vacuum
  • 328193 06/25/10 16:25 benesch hall A target vacuum
  • 328190 06/25/10 16:17 folts checklist/tech
  • 328188 06/25/10 15:23 Markowitz Target to BeO
  • 328179 06/25/10 14:10 Maruyama / P HV trip problem on S2m R isolated and fixed
  • 328174 06/25/10 13:25 bogdanw HV channel for S2m Right PMT for coutner #7
  • 328161 06/25/10 11:45 S. Abrahamya VDC
  • 328160 06/25/10 11:43 A. Camsonne Bad HV channel on left arm slot 12 channel
  • Thursday

  • 328144 06/24/10 23:58 Partridge / Cabling of preshower and S2
  • 328143 06/24/10 23:55 A. Shahinyan L-VDC status
  • 328142 06/24/10 21:51 Ole Hansen USB monitor?
  • 328141 06/24/10 21:41 V. Sulkosky VDC Databases Updated with APEX Optics Matrix
  • 328140 06/24/10 18:25 R. Michaels L-HRS gates restored
  • 328138 06/24/10 18:20 R. Michaels DAQ notes, changes
  • 328124 06/24/10 17:14 R. Michaels story of broken laptop
  • 328123 06/24/10 16:27 LeRose Cycling for Q2/3 magnets
  • 328117 06/24/10 13:32 LeRose Cycling for Q2/3 magnets
  • 328116 06/24/10 12:45 A. Camsonne APEX Left HRS status
  • 328114 06/24/10 12:15 A. Camsonne Left HRS stack back in - Cabling work can start
  • 328110 06/24/10 10:05 A. Camsonne APEX experiment status
  • 328108 06/24/10 03:55 S.Abrahamyan L-arm trigger
  • 328104 06/24/10 00:33 A. Camsonne VDC readback not working
  • 328103 06/24/10 00:21 A. Camsonne [APEX] Experiment status
  • Wednesday

  • 328092 06/23/10 20:10 Takashi/Cliv Cerenkov Cables
  • 328091 06/23/10 20:07 A. Camsonne Temporary laptop in Left Arm
  • 328090 06/23/10 20:06 A. Camsonne S0 powered up
  • 328080 06/23/10 17:33 R. Michaels HOWTO for DAQ
  • 328078 06/23/10 17:12 LeRose Left Q2 Cycling
  • 328077 06/23/10 17:04 R. Michaels DAQ is ready. Today's mods
  • 328061 06/23/10 14:54 jwexler Changed Slug Number to 50
  • 328060 06/23/10 13:27 R. Michaels raster cable changes
  • 328059 06/23/10 12:51 romanip, JIX Moller DAQ test ended
  • 328058 06/23/10 12:45 romanip FADC VME CPU replacing
  • 328057 06/23/10 12:12 A. Camsonne Correction
  • 328056 06/23/10 12:05 A. Camsonne Today Radiation survey
  • 328055 06/23/10 11:30 ZAFAR AHMED PREx WAC note - run 4754
  • 328054 06/23/10 11:27 J. Zhang Moller FADC crate rebooted again
  • 328053 06/23/10 11:19 J. Zhang Moller FADC CPU board is dead
  • 328052 06/23/10 09:42 ZAFAR AHMED PREx WAC note - run 4742
  • 328051 06/23/10 09:22 J. Zhang Start Moller DAQs test for 1000Hz
  • Tuesday

  • 328050 06/22/10 21:17 ZAFAR AHMED WAC note - run 4738
  • 328049 06/22/10 21:05 ZAFAR AHMED WAC note - run 4737
  • 328048 06/22/10 21:01 Clive/Takash Left Arm VDC work
  • 328047 06/22/10 20:52 ZAFAR AHMED WAC note
  • 328046 06/22/10 18:16 pking Changing the INJ crate to "Hall C mode"
  • 328044 06/22/10 16:30 R. Michaels L-HRS DAQ mods
  • 328040 06/22/10 15:12 R. Michaels bad screen on laptop on L-HRS
  • 328036 06/22/10 14:36 Kiadtisak GEM database updated
  • 328011 06/22/10 13:14 ZAFAR AHMED WAC note - run 4733
  • 327998 06/22/10 12:36 ZAFAR AHMED WAC note - run 4722
  • 327993 06/22/10 12:14 ZAFAR AHMED WAC note - run 4718
  • 327989 06/22/10 11:47 ZAFAR AHMED WAC note - run 4717
  • 327965 06/22/10 10:43 A. Camsonne Hall A in restricted access - No radiation watch needed any more
  • 327956 06/22/10 10:08 M Friend PREX CompSimple runs (late entry)
  • 327945 06/22/10 09:12 A. Camsonne Hall A in restricted access with access doors locked
  • Monday

  • 327912 06/21/10 19:28 A. Camsonne Status for tonight in Hall A : no work - swing and owl shift cancelled
  • 327911 06/21/10 17:15 R. Michaels Expt switchover, backups of online accounts
  • 327910 06/21/10 16:54 A. Camsonne Work in hall stopped
  • 327908 06/21/10 16:45 R. Michaels DAQ modifications: run on adaql2 now
  • 327898 06/21/10 14:20 folts/tech target vacuum
  • 327897 06/21/10 13:58 Abdurahim Compton lasers are off
  • 327896 06/21/10 12:41 A. Camsonne Radcon survey finished : access to Hall is possible within the roped boundaries
  • 327895 06/21/10 12:18 Abdurahim Compton epics reboot
  • 327894 06/21/10 10:05 jmammei losing target vacuum
  • 327893 06/21/10 09:55 jmamei target vacuum
  • 327892 06/21/10 09:50 Fansler Moeller warm up
  • 327889 06/21/10 08:51 A. Camsonne Hall A current status : no access permitted for 3 hours
  • 327888 06/21/10 08:07 jmammei Run Coordinator Daily Summary, June 20 Swing - June 21 Owl
  • 327887 06/21/10 07:54 jwexler Tgt. vacuum for the last two shifts
  • 327886 06/21/10 07:38 Trevor Biela vacuum pressure
  • 327885 06/21/10 07:36 jmammei ramped down magnets - except R Q1
  • 327884 06/21/10 07:09 ZAFAR AHMED Target stirp chart. Beam off. Cooler off. Heater on
  • 327883 06/21/10 07:08 jmammei septum turned off
  • 327880 06/21/10 07:01 ZAFAR AHMED MCC has stopped the beam.
  • 327871 06/21/10 06:17 ZAFAR AHMED Detectors Gas System Interlock Status turned ON(TECH)
  • 327868 06/21/10 05:58 ZAFAR AHMED Beam spike
  • 327863 06/21/10 04:28 ZAFAR AHMED Temperateure chart
  • 327860 06/21/10 03:06 jmammei current readback on R Q1
  • 327859 06/21/10 03:06 Trevor Biela q1 right arm detector
  • 327858 06/21/10 02:57 jmammei tried to reset Right Q1
  • 327856 06/21/10 02:17 ZAFAR AHMED Controleld access
  • 327854 06/21/10 01:50 ZAFAR AHMED Q1 in Right arm is down.
  • 327851 06/21/10 01:25 ZAFAR AHMED Q1 in righ arm went down
  • 327850 06/21/10 01:21 ZAFAR AHMED HACWEB1 alaram.
  • 327849 06/21/10 00:59 ZAFAR AHMED Temperature and radition charts.
  • 327848 06/21/10 00:58 ZAFAR AHMED Target strip chart and GUI
  • 327847 06/21/10 00:56 Mindy dithering output from run 4742 to 4745
  • 327844 06/21/10 00:07 M Friend Compton EDET to garage
  • 327843 06/21/10 00:00 Mindy 06/20/2010 Swing shift Summary
  • Sunday

  • 327842 06/20/10 23:58 M Friend adaql5 crashed, back now
  • 327839 06/20/10 23:50 M Friend iocha11 and iocha22 rebooted
  • 327838 06/20/10 23:48 Mindy beampipe temperature sensor broken
  • 327837 06/20/10 23:35 Mindy update calorimeter's temperature
  • 327836 06/20/10 23:26 M Friend spot++
  • 327832 06/20/10 23:24 M Friend T5 rates at 5nA
  • 327828 06/20/10 22:55 Mindy update the calorimeter's temperature
  • 327826 06/20/10 22:53 M Friend Target vacuum snapshot
  • 327816 06/20/10 21:52 Mindy temporature 5.9 now
  • 327813 06/20/10 21:27 M Friend ROC26 is giving LED messages
  • 327812 06/20/10 21:00 Mindy beam energy strip chart
  • 327809 06/20/10 20:28 Selina Target vacuum snapshot
  • 327808 06/20/10 20:03 M Friend L arm HV GUI wrong
  • 327807 06/20/10 19:59 M Friend spot++
  • 327806 06/20/10 19:58 M Friend T5 rates at 5nA
  • 327800 06/20/10 19:45 M Friend Rebooted HRS DAQs
  • 327797 06/20/10 19:26 M Friend Dithering plots
  • 327796 06/20/10 19:26 Mindy calorimeter temperature & right-arm dipole field update
  • 327795 06/20/10 19:03 M Friend Parity DAQ crashed
  • 327789 06/20/10 18:52 A. Camsonne Compton electron detector runs
  • 327788 06/20/10 17:55 A. Camsonne Compton electron detector more analog traces
  • 327787 06/20/10 17:54 Selina Target vacuum snapshot
  • 327786 06/20/10 17:48 Mindy update right-arm dipole strip chart
  • 327783 06/20/10 17:35 Mindy right-arm dipole field and current strip chart
  • 327780 06/20/10 16:45 M Friend Compton FADC flags to CompSimple flags
  • 327779 06/20/10 16:27 M Friend BPM 12x saturating
  • 327778 06/20/10 16:09 jmammei Run Coordinator Daily Summary, June 19 Swing - June 20 Day
  • 327777 06/20/10 16:01 Mindy Lumi and Detector correlation plots in slug 36 and 37
  • 327774 06/20/10 15:49 jmammei right arm settings
  • 327773 06/20/10 15:22 khandler pressure
  • 327771 06/20/10 15:01 ZAFAR AHMED Target - spot++ - T5
  • 327770 06/20/10 15:00 Jack (Techs) Right Dipole controller corrupted, again
  • 327768 06/20/10 14:58 ZAFARF AHMED Left arm scint. rates.
  • 327759 06/20/10 14:30 khandler pressure
  • 327758 06/20/10 14:01 khandler pressure
  • 327756 06/20/10 13:58 ZAFAR AHMED Flipper and new slug is 40
  • 327755 06/20/10 13:57 apar Halfwave plate OUT, Slug 40
  • 327754 06/20/10 13:56 ZAFAR AHMED IHWP is OUT
  • 327752 06/20/10 13:45 ZAFAR AHMED Channel 23 temperture is 580 C
  • 327751 06/20/10 13:23 Mindy Lumi / Detector Correlation in slug 36 / 37
  • 327750 06/20/10 13:14 ZAFAR AHMED WAC note - run 4714
  • 327747 06/20/10 12:52 ZAFAR AHMED Temerature and radition charts
  • 327746 06/20/10 12:40 A. Camsonne Compton electron detector at position -2mm
  • 327745 06/20/10 12:37 khandler Pressure
  • 327742 06/20/10 12:12 A. Camsonne Right arm magnets after cycling magnets
  • 327741 06/20/10 12:04 ZAFAR AHMED Raster check after recycling of quads
  • 327739 06/20/10 12:01 ZAFAR AHMED Right arm magnet vlues are on the tool GUI
  • 327738 06/20/10 12:00 A. Camsonne Turned on regulators on right arm - HAPPEX run 4732 - right arm data not good for this run
  • 327732 06/20/10 11:29 ZAFAR AHMED We are in controlled access
  • 327731 06/20/10 11:12 khandler pressure
  • 327730 06/20/10 11:12 A. Camsonne network 188 in right arm seems to be down
  • 327729 06/20/10 10:51 jmammei right arm readout ioc won't reboot
  • 327726 06/20/10 10:29 ZAFAR AHMED Fast Feedbac GUI
  • 327725 06/20/10 10:28 ZAFAR AHMED General Tool Screen GUI
  • 327724 06/20/10 10:19 ZAFAR AHMED Dithering
  • 327723 06/20/10 10:11 ZAFAR AHMED Temperature and radition stirp charts
  • 327722 06/20/10 10:08 khandler pressure
  • 327721 06/20/10 10:03 ZAFAR AHMED Right magents feedback and PANGUIN
  • 327719 06/20/10 09:50 ZAFAR AHMED Spot++. Run
  • 327716 06/20/10 09:47 ZAFAR AHMED Rates in upper scintillator lefgt arm.
  • 327714 06/20/10 09:26 ZAFAR AHMED Virtual circuit disconnect warning
  • 327713 06/20/10 09:24 ZAFAR AHMED Right spectrometer magnets control VME rebooted.
  • 327712 06/20/10 09:19 khandler Pressure
  • 327711 06/20/10 09:16 ZAFAR AHMED Right arm values on tools GUI are not shown
  • 327710 06/20/10 09:13 ZAFAR AHMED Right arm magnet vales turned white on tool GUI
  • 327707 06/20/10 08:35 ZAFAR AHMED Temperature and radition strip chart.
  • 327706 06/20/10 08:34 jwexler Detector correlations - det3 seems to correlate, det4 doesn't
  • 327705 06/20/10 08:33 khandler pressure
  • 327703 06/20/10 08:00 jwexler Shift Summary - Owl, 6/20
  • 327695 06/20/10 06:56 jwexler False alarm - Compton fine.
  • 327694 06/20/10 06:52 jwexler Compton EPICS needs to be rebooted.
  • 327685 06/20/10 06:49 jwexler spot++
  • 327683 06/20/10 06:37 jwexler ROC 25 crash.
  • 327682 06/20/10 06:05 jwexler ROC6 crashed (again).
  • 327679 06/20/10 05:31 jwexler Dithering and feedback on. Y-coils may need larger amplitudes.
  • 327678 06/20/10 05:29 jwexler Vacuum stripchart
  • 327677 06/20/10 05:28 jwexler Temp. stripcharts
  • 327674 06/20/10 05:00 jwexler White alarm cleared - LeCroy crates elected to start working again.
  • 327673 06/20/10 04:30 jwexler BCM4 not appearing in data output
  • 327672 06/20/10 04:27 jwexler No connectivity to LeCroy crate via Tools GUI - No Compton
  • 327669 06/20/10 04:10 jwexler spot++
  • 327666 06/20/10 03:29 jwexler Collimator temps.
  • 327665 06/20/10 03:25 jwexler Still no HVs
  • 327662 06/20/10 02:47 jwexler No response from LeCroy crates (white alarm)
  • 327661 06/20/10 02:23 jwexler Hall A Tools
  • 327659 06/20/10 02:21 jwexler Target screen & vacuum stripchart
  • 327655 06/20/10 01:21 jwexler Scintillators OFF, going back to production
  • 327654 06/20/10 01:18 jwexler spot++, T5, left arm only
  • 327653 06/20/10 01:17 jwexler ROC6 reboot
  • 327651 06/20/10 01:08 jwexler T5 rates, beam ON
  • 327650 06/20/10 01:06 jwexler Heater off, cooler on.
  • 327649 06/20/10 01:03 jwexler T5 rates for beam OFF
  • 327647 06/20/10 00:41 jwexler Collimator temps (Downstream ~565 deg)
  • 327646 06/20/10 00:35 jwexler Beamline LeCroy HV tripped.
  • 327643 06/20/10 00:02 Abdurahim Compton strip chart
  • Saturday

  • 327642 06/19/10 23:59 Abdurahim Target snapshots
  • 327641 06/19/10 23:23 ZAFAR AHMED WAC note - run 4694
  • 327631 06/19/10 22:47 ZAFAR AHMED Strip charts
  • 327629 06/19/10 22:24 ZAFAR AHMED New Slug - 39
  • 327628 06/19/10 22:23 apar Halfwave plate IN, Slug 39
  • 327627 06/19/10 22:19 Abdurahim spot++ check
  • 327624 06/19/10 21:50 Abdurahim BCM saturation
  • 327623 06/19/10 21:36 Abdurahim Target pressure & temperature
  • 327622 06/19/10 21:34 Abdurahim Target collimator & Radiation Snapshots
  • 327621 06/19/10 21:21 Abdurahim
  • 327620 06/19/10 21:18 Mindy wrong charge in prompt analysis
  • 327617 06/19/10 21:15 Abdurahim BCM 12x saturation !
  • 327616 06/19/10 20:57 Abdurahim Wrong charge value from prompt
  • 327615 06/19/10 20:36 ZAFAR AHMED Strip chart
  • 327612 06/19/10 20:15 ZAFAR AHMED WAC note
  • 327611 06/19/10 20:06 Abdurahim Target strip chart and gui
  • 327610 06/19/10 20:00 ZAFAR AHMED Beam current is 60 uA
  • 327609 06/19/10 19:57 jmammei frequent beam trips
  • 327608 06/19/10 19:49 ZAFAR AHMED Hall A tools GUI
  • 327607 06/19/10 19:49 ZAFAR AHMED Beam and temperature charts
  • 327606 06/19/10 19:40 ZAFAR AHMED Change of configuration
  • 327600 06/19/10 19:22 Abdurahim Compton HV trip
  • 327596 06/19/10 18:53 Abdurahim Lumi Asymmetry
  • 327595 06/19/10 18:33 ZAFAR AHMED Target strip chart and GUI
  • 327592 06/19/10 18:16 Abdurahim raster check
  • 327589 06/19/10 18:07 ZAFAR AHMED Beam is back
  • 327588 06/19/10 18:03 Abdurahim Compton ROC6 reboot
  • 327587 06/19/10 18:00 ZAFAR AHMED High rates right after beam off
  • 327586 06/19/10 17:55 ZAFAR AHMED Beam is down for next 30 minutes
  • 327584 06/19/10 17:19 ZAFAR AHMED Beam condition for last two hours
  • 327583 06/19/10 17:16 Abdurahim Target vacuum temperature & pressure
  • 327580 06/19/10 16:40 ZAFAR AHMED Beam trips
  • 327579 06/19/10 16:15 jmammei Run Coordinator Daily Summary, June 18 Swing - June 19 Day
  • 327578 06/19/10 16:13 M Friend Compton EDET to garage
  • 327574 06/19/10 16:04 M Friend Day shift summary
  • 327573 06/19/10 15:17 khandler pressure
  • 327570 06/19/10 15:10 M Friend Temp stripchart
  • 327562 06/19/10 14:46 khandler pressure
  • 327560 06/19/10 14:23 M Friend Temp monitors back up
  • 327559 06/19/10 14:19 M Friend VDC HVs on/off
  • 327558 06/19/10 14:18 N Toro spot++ R-HRS
  • 327557 06/19/10 14:17 N Toro spot++ more run info
  • 327554 06/19/10 13:53 N Toro spot++
  • 327552 06/19/10 13:51 khandler pressure
  • 327549 06/19/10 13:44 M Friend Trigger rates
  • 327545 06/19/10 13:06 M Friend Lost readback on temps
  • 327544 06/19/10 12:55 M Friend Right arm DAQ back up
  • 327542 06/19/10 12:46 M Friend Compton EDET to -2mm
  • 327541 06/19/10 12:35 N Toro high Lumi 3
  • 327540 06/19/10 12:25 khandler pressure
  • 327539 06/19/10 12:22 M Friend Temp stripchart
  • 327538 06/19/10 12:20 M Friend heater off/blower on
  • 327536 06/19/10 12:17 jmammei VDCs on during runs 6965 and 27922
  • 327535 06/19/10 12:14 khandler Pb Target 1
  • 327534 06/19/10 12:12 N Toro spot++
  • 327530 06/19/10 12:04 M Friend Septum on
  • 327529 06/19/10 11:57 Jixie Zhang Moller has been refilled
  • 327527 06/19/10 11:23 M Friend Trouble with R-HRS DAQ
  • 327526 06/19/10 11:01 M Friend heater off/cooler on
  • 327525 06/19/10 10:59 M Friend Controlled access
  • 327524 06/19/10 10:58 M Friend Septum off
  • 327522 06/19/10 10:45 khandler happy vacuum
  • 327521 06/19/10 10:45 M Friend Hall A beam energy
  • 327518 06/19/10 09:52 N Toro BPM 12 X saturating
  • 327517 06/19/10 09:38 N Toro detector-BCM correlations
  • 327514 06/19/10 08:20 toron rebooted spectrometer DAQs
  • 327512 06/19/10 08:10 M Friend spot++ on Pb/D #1
  • 327508 06/19/10 08:06 jwexler Shift summary - Owl 6/19
  • 327507 06/19/10 07:31 paschke Large Aq run successful without RHWP
  • 327506 06/19/10 07:24 jwexler Scintillators ON
  • 327497 06/19/10 06:37 Trevor Biela vacuum pressure target controls
  • 327495 06/19/10 06:32 jwexler Hall A Tools - Shift Checklist
  • 327493 06/19/10 06:28 jwexler Downstream collimator almost 550 deg.
  • 327491 06/19/10 05:40 jwexler Changed to IHWP OUT
  • 327490 06/19/10 05:39 apar Halfwave plate OUT, Slug 38
  • 327488 06/19/10 05:32 Trevor Biela vacuum pressure 70 microamps
  • 327486 06/19/10 04:46 M Friend Compton EPICS reboot
  • 327484 06/19/10 04:29 jwexler Test run list for 6/19 owl shift.
  • 327483 06/19/10 04:19 jwexler Collimator temp. chart
  • 327482 06/19/10 04:13 jwexler BPM 12 no longer saturating, no idea why.
  • 327481 06/19/10 04:04 jwexler BPM12 saturating
  • 327478 06/19/10 03:49 jwexler Pedestals changed.
  • 327477 06/19/10 03:35 Trevor Biela vacuum pressure 50 60 microamps
  • 327473 06/19/10 02:09 jwexler Dithering, charge feedback on
  • 327471 06/19/10 02:07 jwexler Low beam cut too high for 50uA.
  • 327470 06/19/10 02:05 jwexler Lumi HVs lowered.
  • 327467 06/19/10 01:52 jwexler Gains reset.
  • 327465 06/19/10 01:35 jwexler Gains off for BPM calibration
  • 327463 06/19/10 01:32 Trevor Biela vacuum pressure 70 microamps
  • 327461 06/19/10 01:27 jwexler Reset PC setpoints
  • 327450 06/19/10 00:35 jwexler PC setpoints changed for large charge asymmetry run.
  • 327449 06/19/10 00:32 jwexler Cannot change RHWP setting for large charge asymmetry run
  • 327448 06/19/10 00:28 Trevor Biela vacuum pressure
  • Friday

  • 327443 06/18/10 23:53 zana Shift Summary Swing shift
  • 327442 06/18/10 23:47 A. Camsonne Compton electron detector analog signal / signal to background measurement
  • 327441 06/18/10 23:46 Mindy BPM 12 saturated
  • 327440 06/18/10 23:36 A. Camsonne Compton electron detector to garage position
  • 327439 06/18/10 23:15 zana vacuum pressure
  • 327438 06/18/10 23:14 zana collimator temperature
  • 327428 06/18/10 22:48 zana Pressure target update
  • 327427 06/18/10 22:45 A. Camsonne Compton electron detector to position -2 mm
  • 327426 06/18/10 22:40 jmammei Compton tuning
  • 327425 06/18/10 22:37 Mindy dithering works fine
  • 327424 06/18/10 22:36 M Friend Compton tune successful!
  • 327423 06/18/10 22:33 Mindy BLumi3 is saturated
  • 327422 06/18/10 22:32 M Friend Compton rates after tune
  • 327421 06/18/10 22:28 jmammei Compton tuning
  • 327420 06/18/10 22:15 Mindy Compton tuning finished... go to Production w/ 70uA now
  • 327418 06/18/10 22:12 zana current ramping at 10uA/min
  • 327417 06/18/10 21:54 zana target to lead diamond #1
  • 327415 06/18/10 21:42 Mindy spot++ --- (2)
  • 327413 06/18/10 21:39 zana BPMA X
  • 327412 06/18/10 21:34 Mindy spot++ for beam recovery
  • 327409 06/18/10 21:29 zana Fast beam recovery procedure
  • 327406 06/18/10 21:27 zana target to empty
  • 327405 06/18/10 21:19 A. Camsonne Compton electron detector to garage position
  • 327402 06/18/10 20:33 zana target vacuum and temperature update
  • 327399 06/18/10 19:49 zana temperature and radiation in the collimator
  • 327398 06/18/10 19:35 M Friend Target vacuum screenshot
  • 327397 06/18/10 19:20 M Friend Parity runs during Compton tuning
  • 327395 06/18/10 19:13 zana target moved to Thick Carbon
  • 327394 06/18/10 19:09 Mindy move to thick C12 for Compton tuning
  • 327391 06/18/10 18:58 Mindy spot++
  • 327388 06/18/10 18:56 zana T5 reading
  • 327381 06/18/10 18:31 R. Michaels re: spot on L-HRS
  • 327380 06/18/10 17:51 zana Collimator temperature update
  • 327379 06/18/10 17:49 zana target temperature and vacuum pressure
  • 327377 06/18/10 17:42 M Friend spot++ on Pb/D #1
  • 327373 06/18/10 17:09 jwexler Lowered Right A_T HV
  • 327372 06/18/10 17:07 M Friend Target vacuum snapshot
  • 327371 06/18/10 17:04 jmammei Run Coordinator Daily Summary, June 17 Swing - June 18 Day
  • 327368 06/18/10 16:34 M Friend Day shift summary
  • 327365 06/18/10 16:16 jmammei BPM12 gain too high 4671
  • 327363 06/18/10 16:14 zana compton
  • 327361 06/18/10 16:11 zana temperature readings during the ramp up to 70 uA
  • 327359 06/18/10 16:07 jmammei target vacuum during ramp-up
  • 327356 06/18/10 15:47 jmammei MCC had trouble going above 60uA
  • 327355 06/18/10 15:34 khandler vacuum
  • 327354 06/18/10 15:33 zana temperature Collimator update ramp to 60uA
  • 327353 06/18/10 15:33 jmammei charge feedback on run 4669
  • 327352 06/18/10 15:26 paschke run list during adc gain / pedestal change
  • 327351 06/18/10 15:25 mercado Changed curmon cut
  • 327350 06/18/10 15:14 zana temperatures current at 55uA
  • 327347 06/18/10 15:08 mercado Good Production Run!
  • 327346 06/18/10 15:05 mercado Made Pedestal Changes for BCM calibration
  • 327343 06/18/10 14:49 mercado Changed Gains in CRL for BCMs
  • 327332 06/18/10 14:43 M Friend Beam current to 55uA
  • 327330 06/18/10 14:42 M Friend Compton HV to 1800V
  • 327327 06/18/10 14:37 M Friend Temp stripchart at 60uA
  • 327320 06/18/10 14:16 M Friend Runbird not working
  • 327316 06/18/10 13:53 A. Camsonne Eletron detector analog signal
  • 327315 06/18/10 13:37 M Friend Temp stripchart
  • 327312 06/18/10 13:25 khandler pressure
  • 327311 06/18/10 13:21 A. Camsonne Compton electron detector in position -2mm
  • 327310 06/18/10 13:09 M Friend Compton EDET to garage
  • 327309 06/18/10 12:51 M Friend Compton EDET to -2mm
  • 327308 06/18/10 12:40 M Friend Temperature stripchart
  • 327305 06/18/10 12:10 M Friend Q2 runs
  • 327304 06/18/10 12:09 M Friend heater off/cooler on
  • 327303 06/18/10 11:58 M Friend Lumi 7 signal low
  • 327302 06/18/10 11:54 M Friend Right arm integrating signal
  • 327301 06/18/10 11:50 khandler Target Pressure
  • 327298 06/18/10 11:45 M Friend Septum on (late entry)
  • 327295 06/18/10 11:32 M Friend VDC HVs off
  • 327291 06/18/10 11:24 M Friend VDC HVs on
  • 327286 06/18/10 11:19 M Friend spot++ on Pb/D #1
  • 327278 06/18/10 11:14 M Friend Trigger rates
  • 327277 06/18/10 11:13 jmammei fire alram during sieve investigation
  • 327276 06/18/10 11:12 khandler Pb Diamond 1
  • 327275 06/18/10 11:09 M Friend spot++
  • 327271 06/18/10 11:06 M Friend Trigger rates
  • 327270 06/18/10 11:03 khandler BeO Target
  • 327269 06/18/10 11:00 khandler Empty Target
  • 327268 06/18/10 10:58 M Friend spot++
  • 327265 06/18/10 10:42 M Friend Septum off
  • 327264 06/18/10 10:39 jmammei RHRS sieve was partially in
  • 327263 06/18/10 10:38 M Friend Beamline HV crate rebooted
  • 327262 06/18/10 10:35 jmammei called for escorted access to check the sieve
  • 327261 06/18/10 10:29 jmammei entries erlated to sieve
  • 327260 06/18/10 10:05 A. Camsonne Focal plane distribution - fine binning - t0 not good
  • 327259 06/18/10 09:56 A. Camsonne phi focal plane vs focal plane x run 6957
  • 327258 06/18/10 09:46 A. Camsonne Sieve plots in log scale 6957 and 6840
  • 327257 06/18/10 09:36 A. Camsonne Projection to sieve run 6840 vs run
  • 327256 06/18/10 09:21 A. Camsonne Focal plane X run 6957 vs 6840
  • 327249 06/18/10 08:43 M Friend More spot++
  • 327247 06/18/10 08:40 M Friend Target ion chamber trips
  • 327245 06/18/10 08:35 M Friend spot++
  • 327243 06/18/10 08:32 M Friend Trigger rates
  • 327241 06/18/10 08:12 khandler BeO Target
  • 327239 06/18/10 08:09 ZAFAR AHMED Owl shift summary
  • 327238 06/18/10 08:07 khandler Spot++
  • 327233 06/18/10 07:54 Trevor Biela vacuum pressure
  • 327230 06/18/10 07:26 Fansler Left Diple NMR
  • 327229 06/18/10 07:03 ZAFAR AHMED Beam update
  • 327228 06/18/10 06:58 Trevor Biela vacuum pressure
  • 327218 06/18/10 06:12 ZAFAR AHMED Field sighn check
  • 327217 06/18/10 06:06 Trevor Biela vacuum pressure
  • 327216 06/18/10 06:06 R. Michaels Please check Quad polarities !
  • 327214 06/18/10 05:41 ZAFAR AHMED Beam off. Cooler off. Heater on.
  • 327212 06/18/10 05:33 R. Michaels Ramping to 70 uA
  • 327211 06/18/10 05:02 R. Michaels Resume 70 uA with 1-arm
  • 327210 06/18/10 04:59 R. Michaels magnet screen
  • 327209 06/18/10 04:56 R. Michaels spot_L and _R
  • 327208 06/18/10 04:56 Trevor Biela vacuum pressure
  • 327206 06/18/10 04:52 jwexler VDCs, scintillators OFF
  • 327205 06/18/10 04:50 R. Michaels septum scan
  • 327200 06/18/10 04:39 R. Michaels transverse ok, acceptance check
  • 327199 06/18/10 04:33 R. Michaels investigating R-HRS problem (low rate)
  • 327189 06/18/10 04:14 jwexler Beamline LeCroy crate tripped - unable to restore
  • 327188 06/18/10 04:12 jwexler VDCs, scintillators ON
  • 327182 06/18/10 04:01 Trevor Biela vacuum pressure
  • 327181 06/18/10 03:51 R. Michaels cycling Q2, Q3 on R-HRS
  • 327177 06/18/10 03:07 ZAFAR AHMED Heater is off. Cooler is on.
  • 327176 06/18/10 03:04 ZAFAR AHMED Change in raster 23x9.5(MCC)
  • 327174 06/18/10 02:58 Trevor Biela Target temp above 30 kelvin
  • 327172 06/18/10 02:55 Bellini Beam Raster LeadDiamond Target 1
  • 327169 06/18/10 02:25 ZAFAR AHMED nA at Lead Diamond Target 1
  • 327167 06/18/10 02:19 Trevor Biela Vacuum Pressure
  • 327166 06/18/10 02:18 ZAFARA AHMED nA at BeO
  • 327165 06/18/10 02:08 ZAFAR AHMED Hall tools gui
  • 327164 06/18/10 01:58 Bellini Raster size
  • 327162 06/18/10 01:54 ZAFAR AHMED Change in raster
  • 327158 06/18/10 01:40 ZAFAR AHMED Parity feedback and beam recovery
  • 327157 06/18/10 01:39 mercado Changed curmon cut for 70uA
  • 327152 06/18/10 01:22 Trevor Biela vacuum pressure
  • 327149 06/18/10 01:11 ZAFAR AHMED Importatn parameters
  • 327146 06/18/10 01:07 R. Michaels set lower conv. gains on ADCs
  • 327145 06/18/10 00:45 ZAFAR AHMED Wein flip
  • 327144 06/18/10 00:44 R. Michaels L-HRS dipole issue
  • 327143 06/18/10 00:24 Trevor Biela vacuum pressure strip chart
  • 327141 06/18/10 00:06 zana Swing Shift Summary
  • Thursday

  • 327139 06/17/10 23:40 P Carter spot++
  • 327136 06/17/10 23:28 zana MCC
  • 327135 06/17/10 23:20 zana upgrade from MCC
  • 327134 06/17/10 23:03 zana update vacuum temperature
  • 327133 06/17/10 22:49 zana vacuum pressure jump
  • 327125 06/17/10 22:42 zana left Dipole Field
  • 327120 06/17/10 21:27 zana update on target temperature
  • 327119 06/17/10 21:12 Jack (Techs) Right Dipole status
  • 327118 06/17/10 21:03 zana target temperature update
  • 327113 06/17/10 20:15 zana temperature target
  • 327112 06/17/10 20:04 zana Update from MCC
  • 327111 06/17/10 19:44 zana Magnets ramping up
  • 327109 06/17/10 19:39 zana temperature
  • 327107 06/17/10 19:14 jpchen target cool-down completed
  • 327106 06/17/10 19:01 zana vacuum cool down update
  • 327105 06/17/10 18:58 jpchen switched from warm return to cold return
  • 327104 06/17/10 18:53 zana 70 uA MCC test
  • 327101 06/17/10 18:15 jpchen cool-down progress and cryo ioc problem
  • 327100 06/17/10 18:06 zana connection problem for cryo supply
  • 327099 06/17/10 18:05 jmammei call Bob to establish 70uA
  • 327098 06/17/10 17:55 zana target vacuum update during cooling
  • 327097 06/17/10 17:53 jmammei Run plan and lists for over night
  • 327094 06/17/10 16:54 hsmith Right Dipole
  • 327093 06/17/10 16:53 zana vacuum pressure progress during cool down
  • 327092 06/17/10 16:51 jmammei Run Coordinator Daily Summary, June 16 Swing - June 17 Day
  • 327090 06/17/10 16:42 jpchen Starting Hall A 15K target cool-down procedure
  • 327088 06/17/10 16:36 jmammei target vaccum before cool-down
  • 327087 06/17/10 16:35 folts/tech target vacuum
  • 327085 06/17/10 15:53 jwexler Shift summary - Day 6/16
  • 327082 06/17/10 15:50 jmammei N2 ice flashed off!
  • 327081 06/17/10 15:45 jwexler Target moving to empty
  • 327069 06/17/10 13:49 jwexler Heater on, cooler off.
  • 327068 06/17/10 13:24 jwexler Ramping down all magnets.
  • 327067 06/17/10 13:21 J. Zhang Moller magnet has been refilled
  • 327064 06/17/10 12:46 jpchen warm-up procedure completed. In warm return now.
  • 327063 06/17/10 12:25 Higinbotham Target Warm Up
  • 327062 06/17/10 12:19 jpchen target warm up started.
  • 327061 06/17/10 12:12 A. Camsonne 3 rd channel analog electron detector
  • 327060 06/17/10 12:10 jwexler Target being warmed up
  • 327059 06/17/10 11:55 A. Camsonne compton electron detector analog board cabling
  • 327056 06/17/10 10:58 jwexler Hall will be in controlled access until 5PM
  • 327053 06/17/10 09:43 jwexler Right Q1 back
  • 327052 06/17/10 09:38 jwexler Right Q1 tripped
  • 327049 06/17/10 09:03 jwexler Going to controlled access - full sweep in process.
  • 327046 06/17/10 08:40 jwexler Beam energy chart
  • 327043 06/17/10 08:02 R. Michaels owl shift summary
  • 327042 06/17/10 07:17 R. Michaels move to empty target
  • 327040 06/17/10 07:16 R. Michaels collim. temps (update)
  • 327039 06/17/10 07:12 R. Michaels vacuum red alarm (momentary)
  • 327029 06/17/10 06:43 R. Michaels tungsten temperatures / beam tune
  • 327025 06/17/10 06:07 R. Michaels Compton: too many HV trips
  • 327023 06/17/10 05:39 R. Michaels tools screen
  • 327022 06/17/10 05:38 R. Michaels target screens
  • 327019 06/17/10 04:31 R. Michaels det3,4 signals
  • 327018 06/17/10 04:20 R. Michaels Compton HV to 1850
  • 327017 06/17/10 04:16 R. Michaels Beam ramp up. vacuum.
  • 327016 06/17/10 04:12 R. Michaels start Compton run 23175, HV stable
  • 327015 06/17/10 04:07 R. Michaels restart parity DAQ, start production
  • 327011 06/17/10 03:45 R. Michaels heater off, cooler on
  • 327010 06/17/10 03:40 R. Michaels Compton HV problems
  • 327008 06/17/10 03:35 R. Michaels Pb/D tgt 3 density
  • 327002 06/17/10 03:24 R. Michaels move to Pb/D tgt 3
  • 327001 06/17/10 03:22 R. Michaels turned on septum
  • 327000 06/17/10 03:09 R. Michaels putting in BeO target
  • 326999 06/17/10 03:07 R. Michaels spot++
  • 326995 06/17/10 01:53 R. Michaels beam status: only 5 uA for now
  • 326990 06/17/10 00:16 R. Michaels re: gas shed (no key)
  • 326984 06/17/10 00:01 V. Sulkosky Swing Shift Summary
  • Wednesday

  • 326983 06/16/10 23:55 V. Sulkosky Tune Beam to Hall A
  • 326982 06/16/10 23:51 jwexler Vacuum plot - past 16 hours.
  • 326981 06/16/10 23:36 V. Sulkosky Compton HV Off
  • 326980 06/16/10 23:34 V. Sulkosky Beam Update (Vacuum Issue)
  • 326979 06/16/10 23:30 V. Sulkosky Gas Shed Pressure/Alarm Handler
  • 326978 06/16/10 23:16 V. Sulkosky TECH IOCHA17 (Gas Shed)
  • 326975 06/16/10 22:53 jwexler Target vacuum & shift checklist (tgt. screen)
  • 326967 06/16/10 22:36 V. Sulkosky PSS - Cut Fiber Optic Remote I/O cables in South Linac Summary
  • 326964 06/16/10 22:14 V. Sulkosky Target Vacuum
  • 326963 06/16/10 22:06 V. Sulkosky Re: Relevant ELOG Entries to FIber Optics Problem
  • 326962 06/16/10 21:56 V. Sulkosky Hall A Being Swept
  • 326956 06/16/10 21:09 V. Sulkosky Hall A Tools
  • 326955 06/16/10 21:04 jwexler Target Vacuum
  • 326954 06/16/10 20:47 V. Sulkosky Hall A PSS
  • 326951 06/16/10 20:18 V. Sulkosky Relevant ELOG Entries to FIber Optics Problem
  • 326950 06/16/10 20:18 V. Sulkosky Relevant ELOG Entries to FIber Optics Problem
  • 326949 06/16/10 20:10 V. Sulkosky Beam Status Update
  • 326948 06/16/10 20:03 jwexler Target Vacuum
  • 326945 06/16/10 19:05 jwexler Target Vacuum
  • 326942 06/16/10 18:08 jwexler Target Vacuum
  • 326939 06/16/10 17:38 V. Sulkosky Septum Turned On
  • 326938 06/16/10 17:07 V. Sulkosky Beam Update
  • 326935 06/16/10 16:47 jwexler Target Vacuum
  • 326934 06/16/10 16:45 M Friend Beamline HV crate rebooted
  • 326933 06/16/10 16:35 jmammei plan for tonight
  • 326932 06/16/10 16:25 jmammei Run Coordinator Daily Summary, June 15 Swing - June 16 Day
  • 326928 06/16/10 15:52 folts target vacuum
  • 326927 06/16/10 15:44 Higinbotham Vacuum Inprovement
  • 326925 06/16/10 15:22 Kiadtisak Database updated
  • 326923 06/16/10 15:01 Jixie Zhang Moller magnet has been filled
  • 326911 06/16/10 13:26 Abdurahim Compton epics reboot
  • 326910 06/16/10 12:51 jmammei beam down until 8pm?
  • 326903 06/16/10 10:50 ZAFAR AHMED Change in beam energy during run 4594
  • 326902 06/16/10 10:21 jmammei vacuum alarms
  • 326901 06/16/10 09:44 paschke Beam energy change in run 4594
  • 326900 06/16/10 09:33 ZAFAR AHMED Change in respose of det 3 and 4
  • 326897 06/16/10 08:37 ellie Run #4594 Screenshots: Part III
  • 326896 06/16/10 08:36 ellie Run #4594 Screenshots: Part II
  • 326895 06/16/10 08:35 ellie Run 4594 (last good run) Screenshots: Part I
  • 326892 06/16/10 08:03 ZAFAR AHMED Septum magnet is off
  • 326890 06/16/10 07:39 Markowitz target vacuum
  • 326888 06/16/10 07:38 Markowitz target to empty
  • 326887 06/16/10 07:38 R. Michaels owl shift summary
  • 326885 06/16/10 07:05 Markowitz BeO target in
  • 326884 06/16/10 07:04 Markowitz target vacuum (heater on)
  • 326874 06/16/10 06:37 R. Michaels cooler off, heater on
  • 326873 06/16/10 06:21 R. Michaels beam gone -- ACC problems
  • 326872 06/16/10 06:07 Markowitz target vacuum
  • 326871 06/16/10 05:56 Markowitz target vacuum
  • 326869 06/16/10 05:56 R. Michaels CODA reboot
  • 326864 06/16/10 04:47 R. Michaels tools screen
  • 326863 06/16/10 04:45 R. Michaels target screens
  • 326860 06/16/10 03:50 Markowitz target vacuum
  • 326858 06/16/10 03:10 R. Michaels mini-recovery procedure
  • 326857 06/16/10 03:08 Markowitz target to PB-D1
  • 326856 06/16/10 03:08 Markowitz Target to BeO
  • 326851 06/16/10 02:45 R. Michaels failed to reset R-HRS dipole (TECH)
  • 326850 06/16/10 02:15 Markowitz septum turned off
  • 326849 06/16/10 02:08 Markowitz lost right dipole
  • 326848 06/16/10 02:07 R. Michaels R-HRS dipole is down (TECH)
  • 326846 06/16/10 01:51 R. Michaels det 1,2 flatlined
  • 326843 06/16/10 01:07 Markowitz target vacuum
  • 326841 06/16/10 00:46 apar Halfwave plate IN, Slug 36
  • Tuesday

  • 326838 06/15/10 23:57 V. Sulkosky Swing Shift Summary
  • 326835 06/15/10 23:37 nmuangma target vacuum
  • 326834 06/15/10 23:33 V. Sulkosky Beamline HV Control Issue
  • 326833 06/15/10 23:22 nmuangma target
  • 326830 06/15/10 23:05 V. Sulkosky Compton StripChart
  • 326827 06/15/10 22:59 N Lubinsky Dithering
  • 326817 06/15/10 22:02 V. Sulkosky Radiation Monitors Stripchart
  • 326816 06/15/10 21:53 nmuangma target vacuum/ beam current
  • 326815 06/15/10 21:35 V. Sulkosky Lowered Current to 40 uA
  • 326813 06/15/10 21:22 nmuangma target vacuum / spike
  • 326811 06/15/10 21:02 nmuangma target vacuum/ spike
  • 326810 06/15/10 20:51 N Lubinsky Dithering
  • 326809 06/15/10 20:36 nmuangma target vacuum
  • 326806 06/15/10 19:56 V. Sulkosky Collimator Temperatures
  • 326803 06/15/10 19:32 jmammei more wishful thinking
  • 326802 06/15/10 19:26 jmammei errata for RC Summary
  • 326801 06/15/10 19:16 nmuangma target vacuum/ bad
  • 326800 06/15/10 19:04 jmammei target vacuum after trips
  • 326799 06/15/10 19:02 nmuangma target vacuum/ stable
  • 326798 06/15/10 19:01 N Lubinsky Dithering
  • 326797 06/15/10 18:58 jmammei runs w/ asymptotic temps
  • 326796 06/15/10 18:51 V. Sulkosky Beamline HV Cycled
  • 326793 06/15/10 18:20 nmuangma beam trip/ target vacuum
  • 326792 06/15/10 18:12 nmuangma target vacuum
  • 326789 06/15/10 17:46 V. Sulkosky Hall A Tools Capture
  • 326788 06/15/10 17:43 N Lubinsky Dithering
  • 326787 06/15/10 17:33 N Lubinsky Dithering
  • 326786 06/15/10 17:26 jmammei temperatures with PD#1 vs. #3
  • 326785 06/15/10 17:20 V. Sulkosky MCC Sending Tune Beam
  • 326782 06/15/10 16:53 V. Sulkosky Cycled Beamline HV Crate
  • 326781 06/15/10 16:52 nmuangma beam trip/ target vacuum
  • 326780 06/15/10 16:39 jmammei Run Coordinator Daily Summary, June 14 Swing - June 15 Day
  • 326779 06/15/10 16:37 A. Camsonne Compton new normalized rate
  • 326778 06/15/10 16:28 nmuangma target vacuum
  • 326777 06/15/10 16:23 jwexler Det. HVs for 40uA beam
  • 326774 06/15/10 16:18 jwexler Det. HVs for 50uA beam.
  • 326772 06/15/10 16:02 ZAFAR AHMED June 15 Day Shift Summary
  • 326770 06/15/10 15:59 Jack (Techs) Right Dipole controller corrupted
  • 326769 06/15/10 15:54 jmammei magnet ramp complete?
  • 326768 06/15/10 15:45 jmammei too much positive thinking
  • 326767 06/15/10 15:42 khandler Vacuum Strip
  • 326766 06/15/10 15:40 jmammei magnet ramp complete
  • 326765 06/15/10 15:26 ellie LHRS ps5 and ps8 changed
  • 326764 06/15/10 15:24 ellie spot++ for target check
  • 326762 06/15/10 15:13 ZAFAR AHMED Beam Recovery
  • 326760 06/15/10 15:10 ellie Changing LHRS ps5, ps8
  • 326759 06/15/10 15:07 khandler Vacuum status - Pb 1 tartget
  • 326758 06/15/10 15:02 ZAFAR AHMED Beam Recovery - nA at BeO
  • 326757 06/15/10 14:49 khandler Vacuum status - Beo target
  • 326756 06/15/10 14:44 ellie spot++
  • 326746 06/15/10 14:39 jmammei Right dipole fixed and ramping
  • 326745 06/15/10 14:34 ZAFAR AHMED Beam recovery stopped
  • 326744 06/15/10 14:28 ZAFAR AHMED Beam recovery
  • 326742 06/15/10 14:07 khandler Vacuum during target move
  • 326741 06/15/10 13:57 ellie LHRS ps8 set to 2
  • 326740 06/15/10 13:57 khandler BEAM RECOVERY, EMPTY TARGET
  • 326739 06/15/10 13:54 jmammei begin beam recovery procedure
  • 326738 06/15/10 13:39 ZAFAR AHMED Beam studies are finished
  • 326737 06/15/10 13:09 Abdurahim Compton Cavity has been restored
  • 326736 06/15/10 12:18 ZAFAR AHMED Right dipole is down
  • 326735 06/15/10 12:15 Higinbotham Right Dipole Problem
  • 326734 06/15/10 12:15 J. Zhang Moller magnet has been refilled
  • 326733 06/15/10 11:14 ZAFAR AHMED Off set in run 4570 (WAC)
  • 326732 06/15/10 10:37 M Friend Compton run during large charge asym
  • 326731 06/15/10 10:27 ZAFAR AHMED Controlled access requested
  • 326730 06/15/10 10:06 jwexler Heater on, cooler off
  • 326728 06/15/10 10:00 ZAFAR AHMED Beam studies starts
  • 326725 06/15/10 09:49 ellie HAPPEX Feedback Monitor Screengrab
  • 326724 06/15/10 09:47 ellie Turning Charge Feedback and Ditherting back ON
  • 326722 06/15/10 09:41 Higinbotham Target Burp
  • 326721 06/15/10 09:37 jmammei parity controls screen
  • 326720 06/15/10 09:36 ellie BCM Asymmetries Snapshot
  • 326719 06/15/10 09:35 ellie Feedback Monitor Capture
  • 326717 06/15/10 09:32 ellie Charge Feedback Off, Dithering is Off
  • 326716 06/15/10 09:30 ellie HAPPEX Feedback Monitor Screenshot
  • 326714 06/15/10 08:59 ellie Dithering Plots Screenshots
  • 326713 06/15/10 08:53 ZAFAR AHMED Temperature strip charts.
  • 326712 06/15/10 08:52 Higinbotham Cooling Fan Off & On
  • 326711 06/15/10 08:51 dalton Problem with Phi_FG calculation or in the archiver
  • 326710 06/15/10 08:49 ZAFAR AHMED Cooler is ON
  • 326709 06/15/10 08:35 ZAFAR AHMED Cooler is off
  • 326707 06/15/10 08:32 A. Camsonne Compton electron detector to garage position
  • 326705 06/15/10 08:21 Higinbotham Target Status
  • 326704 06/15/10 08:09 Abdurahim Compton Cavity has been retuned
  • 326703 06/15/10 08:08 R. Michaels owl shift summary
  • 326701 06/15/10 08:03 R. Michaels dithering fixed
  • 326700 06/15/10 07:53 R. Michaels dithering not working ?
  • 326698 06/15/10 07:17 A. Camsonne EPICS reboot complete - CompSimple run 23098 started
  • 326697 06/15/10 07:15 A. Camsonne Compton electron detector hit pattern run 23098
  • 326696 06/15/10 07:05 R. Michaels rescattering probability
  • 326685 06/15/10 06:17 R. Michaels Vacuum at 40 and 50 uA on Pb
  • 326682 06/15/10 05:56 A. Camsonne Compton EPICS crashed - cavity not locking - EPICS reboot
  • 326681 06/15/10 05:48 A. Camsonne Compton flags changed to compsimple flags
  • 326680 06/15/10 05:47 A. Camsonne Compton electron detector in beam position -2 mm
  • 326679 06/15/10 05:39 R. Michaels vacuum during beam ramp up
  • 326678 06/15/10 05:36 R. Michaels spot++ snapshot of target density
  • 326677 06/15/10 05:30 R. Michaels spot++ run 27893
  • 326673 06/15/10 05:11 R. Michaels restart production
  • 326672 06/15/10 05:09 A. Camsonne GEM HV off for production running
  • 326671 06/15/10 05:04 R. Michaels IHWP to OUT. New slug 35
  • 326670 06/15/10 05:02 apar Halfwave plate OUT, Slug 35
  • 326669 06/15/10 05:01 R. Michaels dipole scan: Down
  • 326665 06/15/10 04:37 R. Michaels vacuum flirting with disaster
  • 326643 06/15/10 03:47 A. Camsonne GEM HV on for low current running
  • 326637 06/15/10 03:14 R. Michaels dipole scan: Up
  • 326616 06/15/10 02:26 R. Michaels reboot HAPPEX DAQ
  • 326612 06/15/10 02:14 R. Michaels vacuum. exciting. not.
  • 326584 06/15/10 01:06 Markowitz target vacuum
  • 326581 06/15/10 00:28 R. Michaels Compton DAQ revived
  • 326579 06/15/10 00:23 R. Michaels riding the wave (vacuum)
  • 326576 06/15/10 00:02 R. Michaels ET problems on Compton
  • Monday

  • 326575 06/14/10 23:56 Or Chen Or Chen
  • 326574 06/14/10 23:42 N Lubinsky Vacuum Pressure
  • 326573 06/14/10 23:25 silwal bpm/bcm pedestals, cal run # 4535
  • 326570 06/14/10 23:00 N Lubinsky Vacuum Pressure
  • 326569 06/14/10 22:51 Amber McCrea Dithering
  • 326561 06/14/10 22:31 silwal dithering/charge feedback turned back on; big charge asym test aborted, MCC cant set RHWP to 2500!!!
  • 326557 06/14/10 22:10 N Lubinsky Vacuum Pressure
  • 326554 06/14/10 21:28 silwal dithering/charge feedback off, calibration/big charge asym run underway
  • 326552 06/14/10 20:58 N Lubinsky Vacuum Pressure
  • 326549 06/14/10 19:50 N Lubinsky Vacuum Pressure
  • 326546 06/14/10 19:07 N Lubinsky Vacuum Pressure
  • 326543 06/14/10 18:25 N Lubinsky Vacuum Pressure instability
  • 326542 06/14/10 18:24 Amber McCrea run 4531
  • 326539 06/14/10 18:13 N Lubinsky Vacuum Pressure
  • 326538 06/14/10 18:03 N Lubinsky Vacuum Pressure
  • 326535 06/14/10 17:24 dalton Wien flip story
  • 326534 06/14/10 17:15 N Lubinsky Vacuum Pressure
  • 326533 06/14/10 17:04 jmammei multiple entries
  • 326532 06/14/10 17:02 jmammei Run Coordinator Daily Summary, June 13 Swing - June 14 Day
  • 326531 06/14/10 17:02 jmammei Run Coordinator Daily Summary, June 13 Swing - June 14 Day
  • 326530 06/14/10 16:59 jmammei Run Coordinator Daily Summary, June 13 Swing - June 14 Day
  • 326529 06/14/10 16:58 Or Chen Requesrt to MCC
  • 326528 06/14/10 16:40 M Friend EDET to garage
  • 326525 06/14/10 16:10 N Lubinsky Vacuum Pressure
  • 326523 06/14/10 16:03 M Friend Compton EPICS reboot
  • 326522 06/14/10 15:58 Markowitz Day shift summary
  • 326520 06/14/10 15:43 Markowitz target vacuum
  • 326518 06/14/10 14:56 Markowitz target vacuum
  • 326515 06/14/10 14:44 Jixie Zhang Run Coordinator Summary (6/7 - 6/14)
  • 326514 06/14/10 14:44 Jixie Zhang Run Coordinator Summary (6/7 - 6/14)
  • 326506 06/14/10 14:07 ellie Run 4524: Wide A_T
  • 326501 06/14/10 13:01 mercado Summary of Recent Slugs 156-163
  • 326500 06/14/10 13:00 Markowitz yellow target alarm
  • 326499 06/14/10 12:52 Markowitz target vacuum a bit noisy
  • 326498 06/14/10 12:41 paschke run 4521: no problem, only small change in beam current
  • 326497 06/14/10 12:21 Markowitz right Q1 trip
  • 326494 06/14/10 12:13 jmammei change in detector count rates
  • 326491 06/14/10 11:52 paschke run 4518: problems due to low bpm12 wires
  • 326490 06/14/10 11:51 adaq temperature strip chart
  • 326489 06/14/10 11:49 ellie Run 4521, Detectors showing double peaks
  • 326488 06/14/10 11:40 Markowitz target vacuum
  • 326487 06/14/10 11:35 paschke run 4517: bpm12 gain failure
  • 326483 06/14/10 10:26 Markowitz target vacuum during ramp
  • 326481 06/14/10 10:00 Markowitz target vacuum
  • 326480 06/14/10 10:00 J. Zhang Run Compton during beam recovery
  • 326478 06/14/10 09:57 J. Zhang ROC6 rebooted for compton DAQ.
  • 326476 06/14/10 09:54 ellie spot++
  • 326475 06/14/10 09:51 ellie Slow counting Take III
  • 326472 06/14/10 09:44 Markowitz target vacuum
  • 326471 06/14/10 09:43 ellie spot++ counting really slowly (continued)
  • 326469 06/14/10 09:39 ellie LHRS run counting really slowly
  • 326468 06/14/10 09:35 Markowitz ran flipper now slug 34
  • 326467 06/14/10 09:34 apar Halfwave plate IN, Slug 34
  • 326466 06/14/10 09:33 adaq Compton end run fail! reboot the DAQ
  • 326463 06/14/10 09:14 fansler Moller temps
  • 326461 06/14/10 09:12 ellie LHRS CODA crashed, rebooting
  • 326458 06/14/10 09:08 Markowitz cooler on, heater off
  • 326457 06/14/10 09:08 ellie ROC 31 was rebooted after DAQ crashed
  • 326456 06/14/10 09:00 ellie Run 4518 Replay all strange
  • 326455 06/14/10 08:58 Markowitz target vacuum
  • 326454 06/14/10 08:58 Markowitz target heater on, cooler off
  • 326453 06/14/10 08:57 ellie Run 4517, BPM a bit wild
  • 326452 06/14/10 08:46 jmammei changed the slugnumber to 33 in DB
  • 326451 06/14/10 08:31 ellie DAQ Crashed, lost connection to ROC31
  • 326450 06/14/10 08:01 ZAFAR AHMED June 14 Owl Shift Summary
  • 326446 06/14/10 06:59 ZAFAR AHMED Hall A tools GUI
  • 326445 06/14/10 06:59 M Friend Target GUI snapshot
  • 326444 06/14/10 06:49 M Friend Target vacuum stripchart
  • 326434 06/14/10 06:09 M Friend EDET to garage
  • 326433 06/14/10 05:59 M Friend EDET to -2mm
  • 326430 06/14/10 05:53 jwexler HVs tweaked
  • 326429 06/14/10 05:50 M Friend Target vacuum snapshot
  • 326426 06/14/10 05:32 ZAFAR AHMED IHWP is OUT
  • 326425 06/14/10 05:31 apar Halfwave plate OUT, Slug 163
  • 326424 06/14/10 05:28 M Friend Target to Pb/D #1
  • 326423 06/14/10 05:28 jwexler Blinded test, increased gain run results
  • 326422 06/14/10 05:25 jwexler PbD#1 vs. Ta signal at 10uA
  • 326421 06/14/10 05:19 M Friend Target to tantalum
  • 326417 06/14/10 05:09 M Friend Target to Pb/D #1
  • 326412 06/14/10 05:01 jwexler Blinded test results - Production HV run (4511)
  • 326411 06/14/10 05:00 jwexler Cooler on, blower off.
  • 326410 06/14/10 04:56 M Friend Target vacuum stripchart
  • 326409 06/14/10 04:48 jwexler Access over
  • 326408 06/14/10 04:42 ZAFAR AHMED Hall is in beam permit.
  • 326407 06/14/10 04:42 M Friend Target to BeO
  • 326406 06/14/10 03:59 ZAFAR AHMED Controlled access
  • 326404 06/14/10 03:44 M Friend Target vacuum snapshot
  • 326401 06/14/10 03:03 P Carter no PANGUIN warnings while blinded
  • 326400 06/14/10 02:43 P Carter target vacuum
  • 326399 06/14/10 02:35 ZAFAR AHMED Blinded test starts
  • 326398 06/14/10 02:35 P Carter ROC3 and ROC1 rebooted
  • 326393 06/14/10 02:29 M Friend Target to Pb/D #1
  • 326392 06/14/10 02:27 P Carter spot++
  • 326389 06/14/10 02:15 jwexler Blinders ON
  • 326388 06/14/10 02:14 ZAFAR AHMED Hall is in pebam permit.
  • 326387 06/14/10 02:11 M Friend Target to BeO
  • 326386 06/14/10 01:35 P Carter target vacuum
  • 326385 06/14/10 00:43 ZAFAR AHMED Controlled Access
  • 326383 06/14/10 00:32 M Friend Target vacuum better with no beam
  • 326382 06/14/10 00:28 M Friend Target vacuum spike
  • 326380 06/14/10 00:12 Kiadtisak GEM Efficiency vs ADC cut (Run 27826)
  • 326379 06/14/10 00:08 Mindy 06/13/10 swing shift summary
  • 326378 06/14/10 00:07 Kiadtisak GEM Efficiency vs ADC cut (Run 27800)
  • 326377 06/14/10 00:03 Kiadtisak GEM Efficiency vs ADC cut (Run 27798)
  • Sunday

  • 326376 06/13/10 23:57 Kiadtisak GEM efficiency vs adc cut (Run 27795)
  • 326373 06/13/10 23:00 Markowitz target vacuum
  • 326372 06/13/10 22:49 rouven Left Q1 down
  • 326371 06/13/10 22:48 Mindy left Q1 down, fixed now
  • 326370 06/13/10 22:47 Mindy check - list update: reverse the polarization sign of left-arm Q1
  • 326360 06/13/10 22:02 Markowitz target vacuum
  • 326357 06/13/10 21:09 Markowitz target vacuum
  • 326355 06/13/10 20:59 Mindy change half-wave plate IN
  • 326354 06/13/10 20:59 apar Halfwave plate IN, Slug 162
  • 326352 06/13/10 20:21 Markowitz target vacuum
  • 326349 06/13/10 19:52 Mindy temperature update
  • 326348 06/13/10 19:50 rouven Dithering update
  • 326347 06/13/10 19:48 rouven Dithering
  • 326346 06/13/10 19:44 Mindy radiation strip chart
  • 326345 06/13/10 19:19 Mindy check - list
  • 326344 06/13/10 19:13 Mindy transverse data taking now!!!
  • 326342 06/13/10 19:10 Jack (Techs) LQ1 current set value displayed as zero
  • 326340 06/13/10 19:07 rouven BPM12 saturating
  • 326339 06/13/10 18:59 Jack (Techs) Make that Right Q2
  • 326338 06/13/10 18:56 Markowitz target vacuum during ramping
  • 326335 06/13/10 18:38 Markowitz target vacuum
  • 326330 06/13/10 18:21 Markowitz target vacuum
  • 326328 06/13/10 18:16 Mindy turn the cooler/heater on/off
  • 326322 06/13/10 17:57 rouven R-HRS spot++
  • 326321 06/13/10 17:55 Markowitz target to Pb-1
  • 326320 06/13/10 17:55 hsmith Right Q3
  • 326315 06/13/10 17:44 Mindy turn the septum on
  • 326314 06/13/10 17:37 J. Zhang Going to recover the beam.
  • 326313 06/13/10 17:32 Markowitz target to empty
  • 326311 06/13/10 17:15 Markowitz vacuum pressure
  • 326310 06/13/10 16:38 J. Zhang iocha16 rebooted
  • 326308 06/13/10 16:06 Ole Hansen Day Shift Summary
  • 326307 06/13/10 15:58 Trevor Biela vacuum pressure
  • 326306 06/13/10 15:05 Ole Hansen Hall status
  • 326305 06/13/10 15:03 Jixie Zhang Moller magnet has been refilled
  • 326304 06/13/10 15:01 Trevor Biela vacuum pressure
  • 326303 06/13/10 15:00 Trevor Biela vacuum pressure beam off
  • 326302 06/13/10 14:59 J. Zhang temperature monitor for moller cryo dead
  • 326294 06/13/10 14:41 Ole Hansen Septum off
  • 326293 06/13/10 13:59 Trevor Biela vacuum pressure beam off
  • 326292 06/13/10 13:10 Ole Hansen O-ring cooler off, heater on
  • 326291 06/13/10 13:08 A. Camsonne Right Q2 is not turning on
  • 326289 06/13/10 12:58 Trevor Biela vacuum chamber beam trip
  • 326288 06/13/10 12:30 Ole Hansen Access to reset R-Q2
  • 326287 06/13/10 12:14 Ole Hansen Right Q1 trip
  • 326285 06/13/10 12:14 dalton Root files linked
  • 326283 06/13/10 12:07 Ole Hansen Right Q2 trip
  • 326282 06/13/10 12:04 Trevor Biela vacuum chamber temperature
  • 326279 06/13/10 11:32 Ole Hansen Intermediary Run Summary
  • 326278 06/13/10 11:06 Ole Hansen IHWP OUT
  • 326277 06/13/10 11:03 Trevor Biela vacuum chamber temperature
  • 326275 06/13/10 10:57 apar Halfwave plate OUT, Slug 161
  • 326273 06/13/10 10:53 O. hansen, J /adaql7/work1 filled up
  • 326272 06/13/10 10:42 khandler DISK FULL adaq7
  • 326271 06/13/10 10:41 Ole Hansen Compton HV trip
  • 326270 06/13/10 10:36 A. Camsonne Compton - Beamline HV tripped
  • 326267 06/13/10 10:01 tbielarski vacuum temperature trip
  • 326264 06/13/10 09:03 Trevor Biela Target vacuum snapshot
  • 326262 06/13/10 08:22 jwexler Back to 50uA. Slow ramp complete
  • 326259 06/13/10 07:59 jwexler Shift Summary - Owl 6/13/2010
  • 326257 06/13/10 07:47 M Friend Target vacuum excursion
  • 326256 06/13/10 07:33 jwexler Flipper run - IHWP IN - Slug 160
  • 326254 06/13/10 07:31 apar Halfwave plate IN, Slug 160
  • 326252 06/13/10 06:45 M Friend Hall A bcm slightly wrong
  • 326250 06/13/10 06:41 jwexler Lumis HV for PbD#1
  • 326249 06/13/10 06:41 jwexler PbD#1 Det. HVs.
  • 326240 06/13/10 06:36 M Friend Target vacuum during beam trips
  • 326239 06/13/10 06:01 M Friend Compton FADC sign check
  • 326236 06/13/10 05:37 M Friend Compton EPICS reboot
  • 326235 06/13/10 05:17 jwexler Target Vacuum - 5:15
  • 326232 06/13/10 04:53 jwexler HV GUI constantly crashing
  • 326231 06/13/10 04:39 M Friend Compton vertical scan
  • 326230 06/13/10 04:29 jwexler Add'l troubleshooting notes for the HV GUI
  • 326229 06/13/10 04:26 M Friend Target vacuum stripchart
  • 326224 06/13/10 03:23 M Friend Target vaccum stripchart
  • 326223 06/13/10 03:07 jwexler Beginning ramp on PbD#1
  • 326219 06/13/10 03:00 jwexler Heater off, blower on
  • 326218 06/13/10 02:57 M Friend scintillators off
  • 326217 06/13/10 02:57 jwexler Target Vacuum, 2:55AM
  • 326216 06/13/10 02:56 M Friend spot++, 6x4 raster, x=0, y=0
  • 326211 06/13/10 02:50 M Friend spot++ with 6x4 raster
  • 326207 06/13/10 02:46 M Friend spot++
  • 326200 06/13/10 02:42 M Friend spot++
  • 326191 06/13/10 02:27 M Friend spot++
  • 326186 06/13/10 02:24 M Friend spot++
  • 326181 06/13/10 02:21 M Friend spot++
  • 326176 06/13/10 02:18 M Friend spot++
  • 326173 06/13/10 02:14 M Friend spot++
  • 326168 06/13/10 02:10 M Friend spot++
  • 326165 06/13/10 02:06 jwexler Rates fine on PbD#1.
  • 326164 06/13/10 02:04 M Friend Target to Pb/D #1
  • 326163 06/13/10 02:03 jwexler Slit/attenuator settings for retuned beam
  • 326162 06/13/10 02:00 M Friend spot++
  • 326154 06/13/10 01:46 M Friend position scan
  • 326153 06/13/10 01:43 jwexler Target Vacuum - 1:40AM
  • 326149 06/13/10 00:52 R. Michaels rates on BeO unstable
  • 326148 06/13/10 00:44 M Friend Right Q1 regulation on
  • 326147 06/13/10 00:37 jwexler Target change: BeO
  • 326144 06/13/10 00:36 R. Michaels establishing Pb/D tgt 1 (part 2)
  • 326136 06/13/10 00:22 R. Michaels putting in Pb/D tgt 1
  • 326130 06/13/10 00:14 jwexler Vacuum- 12AM
  • 326122 06/13/10 00:01 jwexler Shift Summary 20:00-00:00
  • 326121 06/13/10 00:00 R. Michaels putting in BeO.
  • Saturday

  • 326120 06/12/10 23:59 jwexler Transverse Asymmetry PbD#1 Runs are slug 159
  • 326116 06/12/10 23:50 jwexler Scintillators on
  • 326114 06/12/10 23:43 jwexler Changed Target: BeO
  • 326113 06/12/10 23:42 don jones spot++ Left
  • 326112 06/12/10 23:42 R. Michaels position scan.
  • 326109 06/12/10 23:01 don jones vacuum stripchart 11 PM
  • 326108 06/12/10 22:59 jwexler Beam up - Recovering to PbD #1
  • 326100 06/12/10 22:00 J. Zhang Heater on and Coller off
  • 326099 06/12/10 21:57 J. Zhang Target Vacuum
  • 326098 06/12/10 21:06 R. Michaels north linac problem
  • 326096 06/12/10 21:04 R. Michaels R-HRS HV working again
  • 326095 06/12/10 20:55 R. Michaels R-HRS HV totally unusable
  • 326094 06/12/10 20:54 Mindy still slug 157 --- NO change/flip in IHWP
  • 326091 06/12/10 20:38 R. Michaels spot++ and C12 FF
  • 326088 06/12/10 20:13 R. Michaels saturating ADCs
  • 326087 06/12/10 20:01 R. Michaels Carbon A_T running
  • 326086 06/12/10 20:00 R. Michaels adjust cavity gains
  • 326083 06/12/10 19:48 R. Michaels Strange BCMCAV
  • 326082 06/12/10 19:45 J. Zhang Shift Summary 16:00-20:00
  • 326081 06/12/10 19:18 J. Zhang Parity DAQ restarted
  • 326076 06/12/10 18:52 J. Zhang Target vacuum: (Thick C12)
  • 326073 06/12/10 17:53 apar Halfwave plate OUT, Slug 158
  • 326070 06/12/10 16:37 Mindy C12 vacuum = 2.9E-05
  • 326068 06/12/10 16:24 Mindy 06/12/10 Day Shift Summary
  • 326066 06/12/10 16:16 jonesdc screenshot
  • 326064 06/12/10 16:14 dalton IHWP moved out
  • 326062 06/12/10 16:11 Mindy C12 vacuum update : 3.7E-05
  • 326060 06/12/10 15:21 Mindy Comton run starts (transverse asymmetry)
  • 326057 06/12/10 15:08 M Friend Compton rates
  • 326056 06/12/10 15:05 Mindy Compton cavity works fine
  • 326055 06/12/10 15:00 Mindy dithering is on; charge feedback is on
  • 326054 06/12/10 14:53 dalton Set detector HV
  • 326053 06/12/10 14:52 dalton Set Lumi HV
  • 326045 06/12/10 14:36 Mindy Compton run 23042 from now
  • 326044 06/12/10 14:29 Mindy 40uA --> 50uA (C12: 2.6E-05)
  • 326041 06/12/10 14:21 mindy 30uA --> 40uA, C12 vacuum = 2.5E-05
  • 326040 06/12/10 14:13 Mindy from 20uA to 30uA (2.5E-05 now)
  • 326039 06/12/10 14:01 mindy C12 vacuum = 4E-05 (50uA 5 mins ago, 10uA now)
  • 326038 06/12/10 14:00 A. Camsonne Cavity monitor calibration attempt
  • 326036 06/12/10 13:31 Mindy spot++ with 5nA on thick C12 (T5 is on)
  • 326034 06/12/10 13:29 Mindy check T5 rate: 5nA on thick C12
  • 326033 06/12/10 13:24 Mindy turn the left-arm upper scintillator on
  • 326031 06/12/10 13:18 Mindy spot++ with 5uA in thick C12
  • 326029 06/12/10 13:04 jwexler Changed to integrating mode
  • 326028 06/12/10 12:39 jwexler Detector HVs set for integrating mode.
  • 326027 06/12/10 12:38 dalton GEMs turned off
  • 326026 06/12/10 12:32 jwexler Scintillators, VDC off
  • 326025 06/12/10 12:21 R. Michaels re: alignment (good)
  • 326024 06/12/10 12:20 dalton Right Q1 tripped
  • 326023 06/12/10 12:16 Mindy Cryo problem causes IOC1 dowm
  • 326019 06/12/10 12:06 dalton Back to standard tune
  • 326017 06/12/10 11:59 dalton Target vacuum update
  • 326016 06/12/10 11:56 Mindy Moving to Thick Carbon
  • 326015 06/12/10 11:49 jwexler R. Arm alignment
  • 326014 06/12/10 11:47 J. Zhang Moller magnet has been filled
  • 326013 06/12/10 11:42 jwexler L. Arm alignment
  • 326012 06/12/10 11:36 Mindy replay backgound scan runs
  • 326010 06/12/10 11:04 dalton Dipole scan
  • 326007 06/12/10 10:44 dalton Left arm at +15%
  • 326006 06/12/10 10:42 dalton Right arm stable +7%, no NMR lock
  • 326000 06/12/10 10:28 dalton Left arm at +10%
  • 325999 06/12/10 10:25 dalton Vacuum flirting with 4e-5
  • 325997 06/12/10 10:00 dalton Changing strategy again
  • 325994 06/12/10 09:56 dalton Left arm at +7%
  • 325989 06/12/10 09:40 dalton Right arm at +4%
  • 325975 06/12/10 08:34 dalton Changing strategy to choose dipole field
  • 325972 06/12/10 08:33 M Friend GEM peaking time scan
  • 325965 06/12/10 08:25 A. Camsonne Going to 2% on left dipole
  • 325963 06/12/10 08:20 M Friend Q2 runs
  • 325961 06/12/10 08:20 A. Camsonne Going to +1 % dipole setting
  • 325955 06/12/10 08:10 M Friend Owl shift summary
  • 325948 06/12/10 07:55 A. Camsonne Production magnet settings
  • 325946 06/12/10 07:50 M Friend Trigger rates at 50nA
  • 325943 06/12/10 07:49 A. Camsonne Thin lead run at 50 nA
  • 325938 06/12/10 07:31 M Friend Target snapshots
  • 325935 06/12/10 07:10 A. Camsonne Misread the rates
  • 325933 06/12/10 06:52 A. Camsonne Resuming thin lead data taking
  • 325928 06/12/10 06:46 A. Camsonne Scaler rates thin lead ~1nA setting
  • 325927 06/12/10 06:44 P Carter thin lead spot++
  • 325922 06/12/10 06:41 M Friend Target to thin lead
  • 325914 06/12/10 06:37 P Carter spot++ plots
  • 325909 06/12/10 06:31 M Friend Target to BeO
  • 325908 06/12/10 06:27 P Carter changed raster size
  • 325905 06/12/10 06:22 M Friend Raster size reduced
  • 325902 06/12/10 06:19 ZAFAR AHMED Cavity box scan
  • 325896 06/12/10 06:14 A. Camsonne Going to Empty target to check beam position
  • 325895 06/12/10 06:07 M Friend Hall A tools snapshot
  • 325894 06/12/10 06:06 M Friend Target screen snapshot
  • 325893 06/12/10 06:05 M Friend Empty target
  • 325892 06/12/10 06:02 ZAFAR AHMED Cavity pregains at 50 nA
  • 325875 06/12/10 05:33 P Carter laser attenuator and slit for BeO
  • 325874 06/12/10 05:32 A. Camsonne Current set for about 10 nA turning detectors back on
  • 325856 06/12/10 05:01 A. Camsonne Scintillator rates at 0.4 uA
  • 325855 06/12/10 04:48 A. Camsonne VDC HV off for BeO check at 0.5 uA
  • 325854 06/12/10 04:47 A. Camsonne PREX detector off for 0.5 uA rates check
  • 325853 06/12/10 04:42 M Friend Target to BeO
  • 325841 06/12/10 04:20 A. Camsonne Qsq Right arm from run 6886
  • 325840 06/12/10 04:19 ZAFAR AHMED Cavity pregains for low current
  • 325835 06/12/10 04:17 A. Camsonne Quartz signal right arm run 6886 - no signal on lower quartz
  • 325832 06/12/10 04:14 M Friend Target snapshots
  • 325824 06/12/10 04:05 A. Camsonne Qsq Left arm run 27804
  • 325823 06/12/10 04:04 A. Camsonne Qsq Left arm run 27804 Quartz signal
  • 325822 06/12/10 03:53 M Friend Target to empty
  • 325821 06/12/10 03:49 ZAFAR AHMED Cavity Pregains for Production Current Restored
  • 325820 06/12/10 03:46 ZAFAR AHMED Access to reboot the cavity IOC
  • 325819 06/12/10 03:44 M Friend Access done
  • 325816 06/12/10 03:32 M Friend Escorted access
  • 325815 06/12/10 03:27 A. Camsonne Asked for escorted access to reboot cavity monitor IOC
  • 325814 06/12/10 03:15 P Carter thick carbon laser attenuator and slit
  • 325811 06/12/10 03:11 A. Camsonne Resuming QsQ measurement while Ahmed is troubleshooting cavity IOC
  • 325810 06/12/10 03:10 M Friend Target to thick carbon
  • 325807 06/12/10 03:06 P Carter spot++
  • 325806 06/12/10 02:59 ZAFAR AHMED IOC of Cavity monitors is down
  • 325803 06/12/10 02:53 P Carter laser attenuator and slit for thin carbon
  • 325802 06/12/10 02:42 M Friend Target to thin carbon
  • 325796 06/12/10 02:16 P Carter laser attenuator and slit
  • 325795 06/12/10 02:08 M Friend Target to BeO
  • 325794 06/12/10 02:04 P Carter spot++ plots
  • 325793 06/12/10 02:03 M Friend Heater off, blower on
  • 325791 06/12/10 01:56 M Friend Beam recovery
  • 325790 06/12/10 01:21 M Friend GEM ADC spectra
  • 325787 06/12/10 00:43 M Friend Update on beam
  • Friday

  • 325785 06/11/10 23:56 Or Chen shift summery
  • 325784 06/11/10 23:29 Mindy after less than one hour, we will get the beam back
  • 325783 06/11/10 23:28 aniol vacuum graph
  • 325775 06/11/10 21:12 A. Camsonne PREX detector HVs set for counting
  • 325774 06/11/10 21:09 A. Camsonne PREX detectors integrating HVs
  • 325773 06/11/10 21:05 A. Camsonne Q2 measurements to be done again
  • 325772 06/11/10 21:03 aniol target moved to empty
  • 325771 06/11/10 20:47 Or Chen MCC
  • 325769 06/11/10 20:19 Mindy Hall-A General Tool
  • 325768 06/11/10 20:17 Kiadtisak Qsq on thin Pb (Right arm)
  • 325767 06/11/10 20:14 Kiadtisak GEM HV Changed
  • 325763 06/11/10 20:05 A. Camsonne Hall A ion chambers
  • 325762 06/11/10 19:50 A. Camsonne Raster current with physics triggers
  • 325760 06/11/10 19:36 A. Camsonne GEM UV chamber u1 v1, HV increased to 3750V
  • 325753 06/11/10 19:25 Jack (Techs) Right Dipole work
  • 325752 06/11/10 19:19 A. Camsonne Run was 27795
  • 325749 06/11/10 19:12 Kiadtisak GEM hit.adc plot
  • 325746 06/11/10 18:58 A. Camsonne INFN GEM HV lowered to 4050V
  • 325743 06/11/10 18:49 aniol vacuum graph
  • 325742 06/11/10 18:45 adaq Q2 value thin Carbon Right arm
  • 325734 06/11/10 18:29 A. Camsonne ROC1 ok after power cycling
  • 325732 06/11/10 18:25 A. Camsonne Right Arm ROC1 fastbus error
  • 325728 06/11/10 18:22 A. Camsonne No rate on Right arm DAQ - rebooted ROC1
  • 325723 06/11/10 18:11 aniol thin carbon Q2 measurement
  • 325722 06/11/10 18:02 A. Camsonne GEM HV on
  • 325721 06/11/10 18:01 Mindy T5 setting: beam recovery
  • 325720 06/11/10 17:56 aniol beam recovery atten and slits settings for BeO
  • 325719 06/11/10 17:30 Or Chen Raster check
  • 325715 06/11/10 16:46 or chen Hall A General Tool
  • 325713 06/11/10 16:43 A. Camsonne Upper scintillator HV and GEM off for target checkout
  • 325711 06/11/10 16:31 jwexler Detectors changed to counting mode
  • 325709 06/11/10 16:17 A. Camsonne Pedestal for GEM updated from run 27792
  • 325708 06/11/10 16:16 ZAFAR AHMED Cavity Monitor Amplifiers is ON
  • 325707 06/11/10 16:01 Boeglin Shift Summary
  • 325705 06/11/10 15:54 R. Michaels run plan for tonight
  • 325703 06/11/10 15:47 A. Camsonne GEM pedestal run 27792
  • 325701 06/11/10 15:45 A. Camsonne HV on on upper scintillator Left HRS
  • 325700 06/11/10 15:43 A. Camsonne GEM checked out
  • 325688 06/11/10 13:35 Jixie Zhang Moller magnet has been refilled
  • 325687 06/11/10 12:03 R. Michaels heater and vacuum status
  • 325686 06/11/10 11:17 mercado BCM1 missing from data stream in slug 33
  • 325685 06/11/10 09:57 Boeglin HiRad Watch necessary
  • 325684 06/11/10 09:16 Fansler/Ewin Target heater
  • 325683 06/11/10 09:08 Boeglin Moeller magnets turned off
  • 325680 06/11/10 08:27 R. Michaels Prex Aux. Plan (proposed)
  • 325678 06/11/10 07:52 C. Gu Target screen update
  • 325677 06/11/10 07:52 jwexler Shift Summary - Owl 6/11/10
  • 325676 06/11/10 07:23 jwexler Called for controlled access - dipole work
  • 325674 06/11/10 06:57 C. Gu Vaccum graph update
  • 325664 06/11/10 06:04 C. Gu Target Screen Update
  • 325663 06/11/10 06:02 C. Gu Vaccum graph update
  • 325660 06/11/10 05:24 jwexler Shift checklist - Hall A Tools
  • 325659 06/11/10 05:16 jwexler Gas alarm panel shows hydrogen both low and high (TECH)
  • 325658 06/11/10 04:59 C. Gu Vaccum graph update
  • 325654 06/11/10 04:03 C. Gu Vaccum graph update
  • 325650 06/11/10 03:03 C. Gu Vaccum graph update
  • 325649 06/11/10 02:24 jwexler Cooler draws current
  • 325647 06/11/10 02:20 jwexler Minor target vacuum alarm tripped
  • 325646 06/11/10 02:12 M Friend CompSimple runs
  • 325644 06/11/10 02:05 C. Gu Vaccum graph
  • 325643 06/11/10 02:03 C. Gu Target move to Pb/D #3
  • 325642 06/11/10 02:03 P Carter spot++ plots
  • 325638 06/11/10 01:59 jwexler Septum on
  • 325633 06/11/10 01:53 jwexler MCC sending tune beam, beginning short beam recovery.
  • 325632 06/11/10 01:49 C. Gu Target move to Empty
  • 325631 06/11/10 01:48 jwexler Heater seems to draw no current(?)
  • 325630 06/11/10 01:11 jwexler MCC says they're waiting on an expert - turning on the heater.
  • 325628 06/11/10 01:01 C. Gu Target Screen Update
  • 325627 06/11/10 00:54 jwexler MCC requested access
  • 325623 06/11/10 00:05 ZAFAR AHMED June 10 Swing Shift Summary
  • 325622 06/11/10 00:04 M Friend CompSimple runs
  • 325621 06/11/10 00:03 A. Camsonne Compton electron detector in garage position
  • Thursday

  • 325620 06/10/10 23:51 ZAFAR AHMED Hall A tools GUI
  • 325619 06/10/10 23:45 A. Camsonne Compton electron detector reduced from 18 to 16 for run 23033
  • 325618 06/10/10 23:40 aniol vacuum graph
  • 325617 06/10/10 23:40 A. Camsonne Compton electron detector in beam position -2 mm
  • 325616 06/10/10 23:34 Mindy beam current is NOT zero
  • 325614 06/10/10 23:30 A. Camsonne Restarted CODA for run 4466
  • 325603 06/10/10 23:06 Kiadtisak Dithering
  • 325595 06/10/10 22:30 A. Camsonn Compton electron detector currents
  • 325594 06/10/10 22:26 M Friend Compton vertical scan
  • 325593 06/10/10 22:20 ZAFAR AHMED HARP scan(Late Entry)
  • 325590 06/10/10 22:04 ZAFAR AHMED Charge Feedback ON
  • 325589 06/10/10 22:02 aniol target charts at beam recovery
  • 325587 06/10/10 21:55 ZAFAR AHMED Heater and Cooler
  • 325585 06/10/10 21:51 Kiadtisak Spot++ Run 27785
  • 325583 06/10/10 21:43 ZAFAR AHMED nA at Lead Diamond # 3
  • 325580 06/10/10 21:35 ZAFAR AHMED nA at BeO
  • 325579 06/10/10 21:23 ZAFAR AHMED septum magnet is ON
  • 325578 06/10/10 21:22 Kiadtisak Spot++ run 27784
  • 325574 06/10/10 20:56 ZAFAR AHMED Beam is back after 5 hours
  • 325573 06/10/10 16:26 Mindy after beam optics: update 4x4 dithering slope history plots and tables
  • 325570 06/10/10 16:16 ZAFAR AHMED Beam Recovery
  • 325568 06/10/10 16:14 aniol target to empty for beam recovery
  • 325566 06/10/10 16:06 Mindy detector asymmetry with 5x5 dithering / regression correction in slug 32
  • 325564 06/10/10 16:05 Boeglinw Shift Summary
  • 325563 06/10/10 15:12 Mindy detector asymmetry with 5x5 dithering / regression correction in slug 33
  • 325555 06/10/10 14:32 J. Zhang Moller magnet has been refilled.
  • 325554 06/10/10 14:14 N Lubinsky Target Screen Capture
  • 325553 06/10/10 14:12 N Lubinsky Tools Capture
  • 325552 06/10/10 14:04 M Friend Compton disk cleanup
  • 325551 06/10/10 13:29 Mindy 5x5 dithering slope table (from slug 32 to slug 33)
  • 325550 06/10/10 13:19 Boeglin HallA survey completed
  • 325549 06/10/10 13:17 Mindy after beam optics: update 5x5 dithering slope history plots
  • 325548 06/10/10 12:32 Boeglin Septum ramped down
  • 325547 06/10/10 11:52 silwal dithering turned OFF
  • 325544 06/10/10 11:36 Boeglin Problem submitting compton runs
  • 325543 06/10/10 11:35 mercado Results of pulse beam study.
  • 325542 06/10/10 11:34 silwal dithering amplitude increased for the last y-coil magnet
  • 325541 06/10/10 11:34 khandler lead target
  • 325539 06/10/10 09:58 N Lubinsky Dithering and Feedback Monitor
  • 325538 06/10/10 09:34 Boeglin Beam off for Machine Optics Studies
  • 325537 06/10/10 09:09 Boeglin Cooler Off/Heater On
  • 325536 06/10/10 09:07 khandler MCC TARGET
  • 325532 06/10/10 08:03 N Lubinsky Dithering and Feedback Monitor
  • 325531 06/10/10 08:01 jwexler Owl shift summary 06/10/10
  • 325529 06/10/10 07:57 C. Gu Target Screen Update
  • 325528 06/10/10 07:25 benesch dithering amplitude
  • 325516 06/10/10 05:48 P Carter parity feedback and dithering plots
  • 325515 06/10/10 05:47 C. Gu Shift Checklist
  • 325514 06/10/10 05:47 jwexler Temp. Stripcharts
  • 325513 06/10/10 05:46 jwexler Shift checklist - Hall A tools
  • 325508 06/10/10 04:43 jwexler Compton ROC6 crashed.
  • 325499 06/10/10 01:20 P Carter parity feedback and dithering plots
  • 325498 06/10/10 00:38 jwexler Turned on dithering
  • 325495 06/10/10 00:19 jwexler Right arm scintillators turned off
  • 325494 06/10/10 00:03 C. Gu Target Screen Update
  • 325493 06/10/10 00:02 jwexler Prompt output shows wrong total charge/beam current.
  • 325491 06/10/10 00:00 ZAFAR AHMED June 9 Swing Shift Summary
  • Wednesday

  • 325488 06/09/10 22:52 Kiadtisak high charge in one single run
  • 325487 06/09/10 22:51 mercado Backed out of pulse beam test
  • 325470 06/09/10 21:42 mercado Setting up for pulse beam test
  • 325468 06/09/10 21:16 aniol target vacuum
  • 325467 06/09/10 21:14 ZAFAR AHMED Hall A Tools Screen GUI
  • 325465 06/09/10 20:50 ZAFAR AHMED I run the flipper for slug 33
  • 325464 06/09/10 20:49 apar Halfwave plate IN, Slug 33
  • 325462 06/09/10 20:43 M Friend Compton background rates
  • 325460 06/09/10 20:27 ZAFAR AHMED Heate is off - Cooler is on
  • 325459 06/09/10 20:21 Mindy No Compton Rate ; Please call Megan
  • 325458 06/09/10 20:13 Kiadtisak Left Spot++ Run 27779
  • 325457 06/09/10 20:11 ZAFAR AHMED Target 3 is a bit damaged
  • 325454 06/09/10 20:04 Kiadtisak Left Spot++ Run 27778
  • 325451 06/09/10 20:01 ZAFAR AHMED nA at Lead Diamond # 3
  • 325450 06/09/10 19:50 ZAFAR AHMED nA at BeO
  • 325449 06/09/10 19:35 Kiadtisak Left Spot++
  • 325446 06/09/10 19:25 Kiadtisak Left spot++
  • 325443 06/09/10 19:17 Kiadtisak Left Spot++
  • 325441 06/09/10 19:11 Kiadtisak Left Spot++
  • 325438 06/09/10 18:12 J. Zhang beam tuning
  • 325437 06/09/10 18:04 ZAFAR AHMED MCC is checking the old tune
  • 325436 06/09/10 18:03 Jack (Techs) Right Dipole
  • 325434 06/09/10 17:57 A. Camsonne rebooted ROC4 from portserver
  • 325432 06/09/10 17:47 Kiadtisak Spot R
  • 325424 06/09/10 17:30 ZAFAR AHMED 5 uA at empty target - Beam recovery
  • 325423 06/09/10 17:19 ZAFAR AHMED Septum magnet is ON
  • 325420 06/09/10 16:27 Jixie Zhang Slug 33 started, HWP in
  • 325418 06/09/10 16:10 R. Lindgren Shift Summary
  • 325413 06/09/10 15:06 silwal Moeller variables added to the halog run epics start/end scripts
  • 325399 06/09/10 13:17 Lindgren Rad watch
  • 325398 06/09/10 13:01 M Friend Compton cptps2 rebooted
  • 325393 06/09/10 10:37 M Friend Compton EPICS reboot problem
  • 325392 06/09/10 10:32 Jixie Zhang Moller solenoid refilled
  • 325391 06/09/10 10:11 lubinsky high radiation watch
  • 325390 06/09/10 10:09 M Friend Bad looking Compton stripchart means bad data
  • 325388 06/09/10 10:00 M Friend Comtpon EPICS reboot
  • 325387 06/09/10 09:49 Jixie Zhang Call up for rad watcher
  • 325386 06/09/10 09:35 J. Zhang Hall A open for control access
  • 325385 06/09/10 09:33 Lindgren Rad Watch
  • 325384 06/09/10 09:09 Lindgren High Rad Watch
  • 325383 06/09/10 09:08 R.Lindgren Target vacuum
  • 325381 06/09/10 08:32 R. Lindgren RHRS Down
  • 325380 06/09/10 08:01 Mindy 06/09/10 shift summary
  • 325379 06/09/10 07:53 C . Gu Target Screen update
  • 325378 06/09/10 07:45 Mindy no Compton rate
  • 325377 06/09/10 07:25 Mindy right arm dipole off
  • 325369 06/09/10 05:39 Mindy turn the raster off before the harp scan
  • 325368 06/09/10 05:37 Mindy harp scan starts
  • 325367 06/09/10 05:32 C. Gu Target Screen update
  • 325366 06/09/10 05:29 Mindy turn the septum off
  • 325365 06/09/10 05:29 C. Gu Target move to Empty
  • 325364 06/09/10 05:28 Mindy turn the beam off
  • 325362 06/09/10 05:14 C. Gu Dithering Online Analysis Screen for Run #4425
  • 325357 06/09/10 02:56 C. Gu CODA restart
  • 325353 06/09/10 02:00 Apar Dithering Online Analysis Screen fro run #4420
  • 325352 06/09/10 01:55 Mindy update target's vacuum ~ 3E-06
  • 325349 06/09/10 01:52 Mindy Compton HV alarm + ROC-6 got stuck
  • 325348 06/09/10 01:22 Mindy not change slug
  • 325345 06/09/10 00:34 Jixie Zhang Run plan for OWL shift of June 9
  • 325344 06/09/10 00:20 C. Gu Target Screen Update
  • 325342 06/09/10 00:12 Mindy Compton HV alarm fixed
  • 325340 06/09/10 00:08 Or Chen Summary
  • Tuesday

  • 325339 06/08/10 23:42 Or Chen Alarm
  • 325325 06/08/10 21:13 aniol vacuum during 40 down
  • 325324 06/08/10 20:50 Jack (Techs) Right Dipole
  • 325322 06/08/10 20:33 J. Zhang Continue data taking with only left HRS
  • 325321 06/08/10 20:10 J. Zhang Right HRS dipole problem
  • 325319 06/08/10 19:44 Kiadtisak Run taken after problem in RHRS
  • 325317 06/08/10 19:22 Or Chen tech dipol Alarm
  • 325316 06/08/10 19:20 or chen tec
  • 325312 06/08/10 18:34 M Friend Compton photon beam position
  • 325310 06/08/10 18:12 Or Chen tools screen
  • 325309 06/08/10 18:11 aniol tgt screens
  • 325308 06/08/10 18:09 M Friend Compton spike in rate
  • 325306 06/08/10 17:25 M Friend Compton background increasing
  • 325301 06/08/10 16:22 jixie Zhang Compton lost connection to ROC6
  • 325300 06/08/10 16:14 Mindy detector asymmetry with 5x5 dithering / regression correction in slug 29
  • 325299 06/08/10 16:12 paschke HCBA injector to hall, run 4406
  • 325298 06/08/10 16:05 R. Lindgren summary
  • 325296 06/08/10 15:43 Mindy detector asymmetry with 5x5 dithering / regression correction in slug 28
  • 325293 06/08/10 15:08 O.Glamadzin Results for Moller large charge asymmetry tests.
  • 325292 06/08/10 14:52 Mindy 5x5 dithering slope table (from slug 28 to slug 31)
  • 325282 06/08/10 13:57 Fansler/Butl Right Q1 Lead flow meters
  • 325279 06/08/10 13:15 M Friend Compton vertical scan
  • 325278 06/08/10 13:08 jwexler Lumi 3 is off
  • 325277 06/08/10 13:05 Lindgren Repairs and checks
  • 325275 06/08/10 13:01 Lindgren Beam start up
  • 325274 06/08/10 12:54 adaq Heater off/Cooler on
  • 325273 06/08/10 12:52 jwexler T5, left arm spot++
  • 325270 06/08/10 12:43 jwexler Slug number changed to 32 for Wien LEFT
  • 325268 06/08/10 12:35 adaq beam recovery raset size
  • 325265 06/08/10 12:09 mcnulty Investigation of counting DAQ quartz det signal problems in Rarm
  • 325264 06/08/10 11:52 Jack (Techs) Septum magnet power supply
  • 325251 06/08/10 11:01 R. Lindgren Target Heater
  • 325248 06/08/10 10:34 Jixie Zhang Moller solemoid has been refilled
  • 325247 06/08/10 10:09 jwexler Late Entry: Dithering fixed
  • 325246 06/08/10 09:47 Mindy dithering amplitude has been changed
  • 325245 06/08/10 09:44 R. Lindgren BigBox
  • 325244 06/08/10 09:20 Jixie Zhang HOW-TO document need to be updated if parameters changed
  • 325243 06/08/10 09:08 Jixie Zhang Control access from 9:00 AM, Tuesday
  • 325242 06/08/10 09:06 R.Lindgren Beam studies
  • 325241 06/08/10 09:05 adaq empty histogram in online monitor
  • 325240 06/08/10 09:04 adaq dithering turned on during 4404
  • 325237 06/08/10 08:41 adaq 4404 w/o dithering
  • 325236 06/08/10 08:39 adaq dithering seems to be off
  • 325234 06/08/10 08:24 R. Lindgren Day Shift start up
  • 325232 06/08/10 08:13 adaq spot_L
  • 325228 06/08/10 08:00 R. Michaels owl shift summary
  • 325227 06/08/10 07:56 R. Michaels reboot parity DAQ
  • 325226 06/08/10 07:44 C. Gu No Beam, Stop run #4402
  • 325225 06/08/10 07:42 R. Michaels Compton reboot, and no beam
  • 325222 06/08/10 06:40 R. Michaels Tgt Flamable Gas Alarm (TECH)
  • 325214 06/08/10 06:15 R. Michaels Hall A tools capture
  • 325213 06/08/10 06:14 R. Michaels target screens
  • 325212 06/08/10 06:05 R. Michaels Compton run stopped 2 min late
  • 325210 06/08/10 06:02 R. Michaels blumi3 HV adjustment
  • 325209 06/08/10 06:00 R. Michaels slug 31 start
  • 325208 06/08/10 05:57 apar Halfwave plate OUT, Slug 31
  • 325200 06/08/10 03:18 R. Michaels HV alarm, beamline crate
  • 325199 06/08/10 03:07 R. Michaels smooth beam delivery
  • 325194 06/08/10 01:12 C. Gu Target Screen Updated
  • 325193 06/08/10 01:00 R. Michaels quality control checks
  • 325190 06/08/10 00:27 R. Michaels Alarm on Q1 Lead flow num2 (TECH)
  • 325188 06/08/10 00:02 zana Shift Summary
  • 325187 06/08/10 00:02 R. Michaels Compton status (good)
  • Monday

  • 325182 06/07/10 22:44 M Friend Compton vertical scan
  • 325172 06/07/10 21:43 zana target vacuum and temperature snapshot
  • 325171 06/07/10 21:40 M Friend Compton background
  • 325170 06/07/10 21:31 Mindy dithering amplitude changed after Moller
  • 325166 06/07/10 21:03 silwal different dithering responses before/after Moller
  • 325165 06/07/10 20:50 Mindy at Rupesh's request: update dithering plots
  • 325164 06/07/10 20:50 M Friend Compton EPICS reboot
  • 325163 06/07/10 20:49 mindi Production run started
  • 325162 06/07/10 20:45 M Friend Results for Compton large charge asymmetry run
  • 325161 06/07/10 20:40 silwal large charge asym run, IHWP IN, -ve asym
  • 325160 06/07/10 20:40 Silwal Parity gui snapshot
  • 325159 06/07/10 20:38 Mindy production: dithering on ; feedback on
  • 325156 06/07/10 20:21 zana vacuum vs beam current
  • 325155 06/07/10 20:15 zana heater turned off and cooler on
  • 325150 06/07/10 19:46 Mindy spot++: ps5=65535; ps8=10
  • 325147 06/07/10 19:43 zana target in lead diamond #3
  • 325146 06/07/10 19:42 J. Zhang Scater rates Pb=#3 , 5nA
  • 325145 06/07/10 19:37 Jixie Zhang Scalers rates BeO 5 nA
  • 325144 06/07/10 19:23 Mindy spot++ in 10uA w/ BeO target
  • 325141 06/07/10 19:16 zana target moved to BeO
  • 325140 06/07/10 19:03 Mindy L-HRS : spot++ (PS8=2;PS5=65535)
  • 325136 06/07/10 18:58 rupesh DAQ problem solved
  • 325133 06/07/10 18:50 O.Glamazdin Moller results
  • 325122 06/07/10 17:59 mindy Parity gui snapshot
  • 325121 06/07/10 17:14 M Friend Compton FADC DAQ to 120Hz
  • 325120 06/07/10 16:10 silwal big charge asym run # 4353, IT 3110, IHWP OUT, asym is +ve
  • 325119 06/07/10 16:08 Boeglin Shift Summary
  • 325117 06/07/10 14:44 M Friend Compton ts crate added to hareboot4
  • 325107 06/07/10 13:48 O.Glamazdin Moller magnet has been filled
  • 325106 06/07/10 13:44 fansler RQ1 Lead Flow
  • 325105 06/07/10 13:23 J. Zhang Escort access to fill the moller
  • 325104 06/07/10 12:53 O.Glamazdin Moller quads power supply is turned OFF
  • 325103 06/07/10 12:50 O.Glamazdin Helicity frequency is 30Hz for Moller
  • 325102 06/07/10 12:45 J. Zhang The dithering was turned off
  • 325100 06/07/10 12:25 M Friend Compton FADC DAQ to 30Hz
  • 325092 06/07/10 12:14 Amber McCrea target change
  • 325090 06/07/10 12:02 silwal dithering x-amplitudes increased to 70 mA
  • 325088 06/07/10 11:51 Boeglin Stopped beam for access
  • 325087 06/07/10 11:50 Boeglin Heater on
  • 325085 06/07/10 11:33 adaq Hall A Tools Screen
  • 325084 06/07/10 11:30 Amber McCrea target screen capture
  • 325083 06/07/10 11:30 Amber McCrea target screen capture
  • 325078 06/07/10 10:08 Boeglin IHWP in
  • 325076 06/07/10 10:03 apar Halfwave plate IN, Slug 30
  • 325074 06/07/10 09:13 boeglin run 4366 very short
  • 325066 06/07/10 08:03 efuchey 6/07/2010 Owl shift summary
  • 325065 06/07/10 07:48 etile charge feedback
  • 325064 06/07/10 07:22 efuchey Beamline HV tripped
  • 325061 06/07/10 07:02 efuchey more about compton rate evolution
  • 325060 06/07/10 06:56 efuchey Compton rates
  • 325059 06/07/10 06:42 etile charge feedback
  • 325058 06/07/10 06:41 etile Dithering BPM/Det
  • 325048 06/07/10 06:25 etile charge feedback
  • 325047 06/07/10 05:40 etile charge feedback
  • 325044 06/07/10 05:03 etile charge feedback
  • 325043 06/07/10 04:56 Sereres J vacuum snapshot
  • 325042 06/07/10 04:55 etile Dithering BPMs/Dets
  • 325041 06/07/10 04:54 efuchey Hall A tools snapshot
  • 325039 06/07/10 04:44 Sereres J target moved
  • 325037 06/07/10 04:36 etile raster size
  • 325033 06/07/10 04:28 A. Camsonne Power cycled network switch in right arm nose beam
  • 325032 06/07/10 04:05 efuchey consequences of IOC crash
  • 325030 06/07/10 04:01 Sereres J Moving to empty target
  • 325029 06/07/10 03:52 A. Camsonne Right arm 188 network down
  • 325028 06/07/10 03:48 efuchey IOCcrash: correction
  • 325027 06/07/10 03:44 efuchey IOCs crash
  • 325026 06/07/10 03:42 etile dithering BPM/Det
  • 325025 06/07/10 03:39 etile charge feedback
  • 325020 06/07/10 02:27 Sereres J Screen captures
  • 325017 06/07/10 01:28 etile BCM asymmetries
  • 325016 06/07/10 01:26 Sereres J target vacuum update
  • 325015 06/07/10 01:25 silwal transmission thruogh the injector/hall bpms/bcms, run # 4356
  • 325013 06/07/10 00:56 silwal dithering current amplitude reduced by 50% in x coils
  • 325010 06/07/10 00:46 efuchey Hall A beam current status
  • 325009 06/07/10 00:44 silwal pedestals for IT 3110
  • 325008 06/07/10 00:41 etile lumi
  • 325007 06/07/10 00:08 A. Camsonne Charge feedback ok
  • 325005 06/07/10 00:03 Mindy 06/06/10 Swing Shift Summary
  • Sunday

  • 325001 06/06/10 23:51 Mindy turn the cooler on; turn the heater off
  • 324999 06/06/10 23:49 A. Camsonne Long target beam recovery procedure complete ramping to 50 uA for production
  • 324997 06/06/10 23:48 A. Camsonne HV scintillators off for production
  • 324995 06/06/10 23:46 A. Camsonne Raster current with T5 on right arm
  • 324990 06/06/10 23:43 A. Camsonne Error message spot++ run 6858
  • 324981 06/06/10 23:35 A. Camsonne Scalers rates Pb #3 ~5nA
  • 324980 06/06/10 23:34 or chen target swich
  • 324979 06/06/10 23:28 Mindy L-HRS spot++
  • 324977 06/06/10 23:26 Mindy L-HRS T5 (long) beam recovery
  • 324974 06/06/10 23:23 Mindy turn the septum on
  • 324973 06/06/10 23:12 Mindy l-HRS spot++ in BeO w/ 10uA (long beam recovery)
  • 324971 06/06/10 23:07 Mindy target to BeO
  • 324970 06/06/10 23:02 Mindy empty target ; target's vacuum ~ 4.7E-07
  • 324969 06/06/10 22:59 Mindy L-HRS spot++ : (long) beam recovery
  • 324967 06/06/10 22:54 Mindy Hall-A general tool : (long) beam recovery
  • 324964 06/06/10 22:37 M Friend Compton orbit lock
  • 324953 06/06/10 22:24 Mindy pedestal test : beam strip chart
  • 324947 06/06/10 21:55 Mindy BPM Gain Off : pedestal test
  • 324943 06/06/10 21:44 A. Camsonne iocse10 fixed
  • 324942 06/06/10 21:40 A. Camsonne Swapped FADC MOLLER VME CPU
  • 324941 06/06/10 21:26 silwal IT changed to 3110, new adc board timing schematic
  • 324935 06/06/10 19:58 A. Camsonne Beam charge asymmetry after restoration of production settings
  • 324934 06/06/10 19:51 M Friend Compton beam charge asymmetry
  • 324933 06/06/10 19:48 Mindy MCC called for the control access in 20 minutes
  • 324932 06/06/10 19:37 A. Camsonne Target BPMs still broken
  • 324931 06/06/10 19:35 Mindy helicity pattern change back to 120hz
  • 324929 06/06/10 19:33 Mindy back out of big charge asym setpoints
  • 324928 06/06/10 19:31 silwal charge feedback on
  • 324927 06/06/10 19:26 M Friend Helicity back to 120Hz
  • 324926 06/06/10 19:11 Mindy change helicity pattern
  • 324925 06/06/10 19:09 silwal big charge asym run # 4344
  • 324924 06/06/10 19:08 M Friend Results for Compton large charge asymmetry run
  • 324921 06/06/10 18:41 Mindy charge asymmetry starts ; ask 50uA
  • 324918 06/06/10 18:36 silwal dithering turned off for big charge asym run
  • 324917 06/06/10 18:36 brads re: Moller FADC VME CPU does not responded
  • 324916 06/06/10 18:26 romanip Moller FADC VME CPU does not responded
  • 324915 06/06/10 18:17 O. Glamazdin Moller has been refilled
  • 324914 06/06/10 18:16 silwal feedback turned off for big charge asym run
  • 324911 06/06/10 17:54 A. Camsonne Access for iocsee10
  • 324910 06/06/10 17:47 Mindy turn the heater on ; turn the cooler off
  • 324909 06/06/10 17:41 Or Chen target vacuum update - 5.5E-07
  • 324908 06/06/10 17:33 Mindy empty target w/ 50uA beam for charge asymmetry measurement in parity DAQ and Moller DAQ
  • 324901 06/06/10 16:59 Mindy MCC called : IOC problem
  • 324900 06/06/10 16:42 or chen production
  • 324899 06/06/10 16:27 Mindy septum off -- control access
  • 324898 06/06/10 16:22 Mindy MCC fixed IOC problem
  • 324897 06/06/10 16:22 A. Camsonne Parity feedback runbird
  • 324893 06/06/10 16:04 Boeglin Shift Summary
  • 324891 06/06/10 15:48 Boeglin Right Q1 is back, affected run 4340
  • 324890 06/06/10 15:40 Boeglin Right Q1 trip
  • 324887 06/06/10 14:57 A. Camsonne Production at 50 uA
  • 324886 06/06/10 14:46 M Friend Compton vertical scan
  • 324877 06/06/10 14:23 Boeglin Switch back to cooling
  • 324876 06/06/10 14:21 Boeglin Switch to heating
  • 324875 06/06/10 14:08 Boeglin Beam off for short time
  • 324869 06/06/10 12:43 Boeglin Hall A Tools screen
  • 324868 06/06/10 12:42 de la codre Target screen capture
  • 324865 06/06/10 12:17 khandler Dithering run 4334
  • 324864 06/06/10 11:46 M Friend Compton rates stripchart
  • 324863 06/06/10 11:46 jwexler Panguin running again
  • 324861 06/06/10 11:34 A. Camsonne Requested 42 uA to check if injector
  • 324859 06/06/10 11:32 M Friend Compton EPICS back up
  • 324858 06/06/10 11:27 A. Camsonne PanGuin broken
  • 324854 06/06/10 11:07 A. Camsonne Injecter BLM trips when going to 50 uA staying around 40 uA while investigating
  • 324852 06/06/10 10:58 M Friend Compton EPICS reboot
  • 324850 06/06/10 10:56 M Friend Compton cavity locking, readback still dead
  • 324847 06/06/10 10:39 khandler spot++ Right Arm
  • 324840 06/06/10 10:33 Boeglinw Cooler on
  • 324837 06/06/10 10:30 A. Camsonne Trigger scintillator HV off
  • 324836 06/06/10 10:29 boeglin spot_R Pb #3
  • 324833 06/06/10 10:27 boeglin spot_L Pb #3
  • 324832 06/06/10 10:27 A. Camsonne Rates on Pb target about 5 nA
  • 324830 06/06/10 10:26 de la codre target moved to lead 3
  • 324827 06/06/10 10:22 boeglinw spot_R BeO
  • 324826 06/06/10 10:22 A. Camsonne Scaler rates with BeO nA
  • 324825 06/06/10 10:20 Boeglin spot_L BeO
  • 324821 06/06/10 09:54 pierre henri back to BeO target
  • 324820 06/06/10 09:52 dalton Source studies last night
  • 324819 06/06/10 09:50 pierre henri change target to lead target 3
  • 324810 06/06/10 09:35 adaq Scintillator trigger on
  • 324806 06/06/10 08:03 ZAFAR AHMED June 6th Owl Shift Summary
  • 324805 06/06/10 07:33 A. Camsonne Compton EPICS reboot
  • 324804 06/06/10 07:00 C. Gu Spot++ Plot
  • 324803 06/06/10 06:59 Sereres J target moved to BeO
  • 324802 06/06/10 06:56 A. Camsonne Beam recovery
  • 324791 06/06/10 05:55 Sereres J target vacuum update
  • 324790 06/06/10 05:39 Sereres J cryo update
  • 324789 06/06/10 02:46 Sereres J Target vacuum update
  • Saturday

  • 324769 06/05/10 23:15 A. Camsonne Argon and Nitrogen cylinder swapped
  • 324764 06/05/10 23:01 Mindy helicity pattern changed
  • 324756 06/05/10 22:25 A. Camsonne [Late entry]Septa magnet ramped down
  • 324755 06/05/10 22:22 A. Camsonne PREX experiment status Saturday June 5th 2010 - No beam for now
  • 324754 06/05/10 21:02 Mindy target vacuum update -- 4.3E-07
  • 324753 06/05/10 20:04 jwexler Backed out of linearity test
  • 324752 06/05/10 19:53 J. Wexler helicity frequency back to 120Hz
  • 324746 06/05/10 19:26 Mindy target's vacuum ~ 4.7e-07
  • 324745 06/05/10 19:18 Mindy beam down for at least a couple of hours
  • 324736 06/05/10 17:55 A. Camsonne Dump Humidity level high
  • 324731 06/05/10 17:02 J. Wexler linearity test : change the helicity frequency
  • 324730 06/05/10 16:59 jixie Moller solenoid has been refilled
  • 324728 06/05/10 16:35 A. Camsonne Update about beam : at best in two hours
  • 324727 06/05/10 16:02 Boeglin Shift summary
  • 324725 06/05/10 15:57 Abdurahim Compton Cavity has been optimized
  • 324724 06/05/10 15:55 M Friend Target to emtpy
  • 324723 06/05/10 15:42 M Friend ROC23 rebooted
  • 324710 06/05/10 13:44 Boeglin Cooler/Heater Status change
  • 324708 06/05/10 13:41 boeglin Beam lost due to 2k cold box trip
  • 324706 06/05/10 13:10 M Friend Target screenshots
  • 324705 06/05/10 13:07 adaq Hall A Tools Screen
  • 324702 06/05/10 12:44 P Carter another online analysis plot
  • 324701 06/05/10 12:43 P Carter online analysis plots
  • 324700 06/05/10 12:25 M Friend Compton vertical scan
  • 324697 06/05/10 11:51 paschke HCBA injector to hall, run 4299, slug 29
  • 324696 06/05/10 11:44 paschke HCBA injector to hall, run 4294, slug 28
  • 324695 06/05/10 11:41 M Friend Compton rates stripchart
  • 324694 06/05/10 11:32 P Carter dither plots
  • 324692 06/05/10 11:19 A. Camsonne Trigger scintillators off
  • 324691 06/05/10 11:15 P Carter another spot++ plot
  • 324681 06/05/10 11:05 PCarter spot++ plots
  • 324674 06/05/10 10:53 M Friend Target to Pb/diamond #3
  • 324673 06/05/10 10:33 P Carter spot++ plots
  • 324671 06/05/10 10:29 Boeglin Cooler on
  • 324670 06/05/10 10:13 M Friend Target snapshot
  • 324669 06/05/10 10:11 P Carter Dithering turned off
  • 324668 06/05/10 10:06 Boeglin Heater on
  • 324667 06/05/10 10:00 dalton Changes to GUI
  • 324666 06/05/10 09:52 A. Camsonne Tune through Compton chicane will affect HAPPEX runs greater than 4304
  • 324664 06/05/10 09:36 A. Camsonne Happex DAQ restarted
  • 324662 06/05/10 09:31 A. Camsonne IOCHA1 rebooted but this did not fix the HARP 4B
  • 324661 06/05/10 09:28 P Carter ROC25 crashed
  • 324659 06/05/10 09:23 adaq BCM calibration constants
  • 324658 06/05/10 09:21 M Friend Target moved to empty
  • 324653 06/05/10 08:25 P Carter "Bad!" quoth the GUI
  • 324651 06/05/10 07:59 benesch response of X and Y bpms to RF indicates dispersion leakage
  • 324650 06/05/10 07:59 jwexler Shift Summary: Owl, 06/05/2010
  • 324649 06/05/10 07:56 C. Gu Target Screen
  • 324645 06/05/10 06:31 R. Barbieri "BAD!" in panguin - IT_qweak related
  • 324637 06/05/10 06:18 jwexler Online analysis reads avg. current = 0. Prompt analysis OK
  • 324635 06/05/10 06:04 jwexler IHWP changed to OUT, SLUG 29
  • 324634 06/05/10 06:02 apar Halfwave plate OUT, Slug 29
  • 324632 06/05/10 05:47 C. Gu Shift summary
  • 324631 06/05/10 05:46 R. Barbieri dither plots
  • 324630 06/05/10 05:45 jwexler Hall A Tools screen (shift summary)
  • 324627 06/05/10 05:32 A. Camsonne HV gui back
  • 324625 06/05/10 05:28 jwexler HV gui up
  • 324624 06/05/10 05:16 A. Camsonne HV gui still down
  • 324623 06/05/10 04:59 A. Camsonne iocha22 up
  • 324620 06/05/10 04:32 jwexler Still no response from HV crates - rebooting again
  • 324619 06/05/10 04:22 R. Barbieri dither plots
  • 324618 06/05/10 04:02 jwexler IT_Qweak drifting above TStable
  • 324615 06/05/10 03:26 jwexler White alarm for HV crates - Rebooted IOCs
  • 324614 06/05/10 03:18 jwexler Beamline Lecroy crate trip
  • 324611 06/05/10 02:28 jwexler Beamline HV crate tripped
  • 324610 06/05/10 02:24 jwexler Beamline thermocouple still flakey
  • 324609 06/05/10 02:18 R. Barbieri production
  • 324602 06/05/10 01:56 R. Barbieri spot++, T5
  • 324599 06/05/10 01:52 jwexler Beam recovery - cooler on, heater off.
  • 324596 06/05/10 01:49 C. GU Target move to Pb/D #3
  • 324594 06/05/10 01:45 C. Gu Target move to BeO
  • 324590 06/05/10 01:39 jwexler Raster size
  • 324589 06/05/10 01:34 R. Barbieri spot++
  • 324588 06/05/10 01:33 jwexler Beam back, beginning recovery
  • 324580 06/05/10 00:04 yawei target screen
  • 324578 06/05/10 00:03 Ole Hansen Swing Shift Summary
  • Friday

  • 324577 06/04/10 23:53 Ole Hansen iochacp2 followup
  • 324576 06/04/10 23:30 Ole Hansen iochacp2 down
  • 324568 06/04/10 21:25 pking ROC31 seems to be working now; changed nsamp to 905
  • 324567 06/04/10 21:08 yawei target moved to empty
  • 324529 06/04/10 17:55 Jixie Zhang Moller solenoid has been refilled
  • 324528 06/04/10 17:51 A. Camsonne Cryo 2K coldbox trip - no beam - 4 hr escalation
  • 324526 06/04/10 17:09 A. Camsonne PREX septum magnet ramped down for access
  • 324525 06/04/10 16:57 Ole Hansen Controlled Access for Moller refill
  • 324524 06/04/10 16:49 Ole Hansen Beam down
  • 324523 06/04/10 16:45 Apar Some more pulsed beam test results
  • 324519 06/04/10 16:09 B. Quinn Day shift summary
  • 324517 06/04/10 16:07 yawei target screen
  • 324516 06/04/10 16:07 mercado Lumi HV cables shifted two channels
  • 324514 06/04/10 15:24 mercado Dither cycle
  • 324513 06/04/10 15:23 B. Quinn Lost communication with Right HV crate
  • 324512 06/04/10 15:13 mercado Lumi 7 saturated... fixed...
  • 324500 06/04/10 13:51 B. Quinn Description of new shiel wall on Compton
  • 324499 06/04/10 13:35 ZAFAR AHMED Access for Moller refill is delayed till 8:00PM
  • 324496 06/04/10 12:38 B. Quinn Hall A tools
  • 324493 06/04/10 12:19 M Friend Compton neutron background rates
  • 324492 06/04/10 12:16 A. Camsonne Access for Moller dewar filling scheduled for 2PM
  • 324491 06/04/10 12:09 A. Camsonne Temperature and pressure fluctuaction maybe related to Fast Feedback issues
  • 324490 06/04/10 12:05 B. Quinn Dithering screen shot
  • 324489 06/04/10 12:00 M Friend Compton neutron background stripchart
  • 324486 06/04/10 11:40 A. Camsonne Compton dipole chicane fixed
  • 324485 06/04/10 11:39 ZAFAR AHMED Target Screen GUI
  • 324484 06/04/10 11:26 B. Quinn Heater off
  • 324483 06/04/10 10:54 B. Quinn cooling profile
  • 324481 06/04/10 10:48 etile raster
  • 324478 06/04/10 10:40 ZAFAR AHMED Tarted is Lead Diamond 3
  • 324477 06/04/10 10:39 silwal vqwk nsamples changed from 956 to 925.
  • 324476 06/04/10 10:37 etile raster
  • 324474 06/04/10 10:23 B. Quinn Change to empty target
  • 324472 06/04/10 10:20 ZAFAR AHMED Target is empty target
  • 324468 06/04/10 09:52 Nilanga Bad sive pattern
  • 324467 06/04/10 09:41 B. Quinn Heater on
  • 324466 06/04/10 09:10 etile Prex config still not working.
  • 324465 06/04/10 09:05 A. Camsonne GSO HV on for Compton background check out from the chicane
  • 324462 06/04/10 09:00 mcnulty Rarm alignment plots from last night
  • 324460 06/04/10 08:56 mercado Lumi HV before/after
  • 324459 06/04/10 08:50 mercado Lumis were off
  • 324458 06/04/10 08:41 jwexler Short signal size comparrisons runs
  • 324457 06/04/10 08:37 mcnulty Quartz ADC dists from last night's alignment runs
  • 324456 06/04/10 08:37 R. Michaels snapshot of temps
  • 324455 06/04/10 08:22 jwexler Charge FFB and dithering ON
  • 324453 06/04/10 08:06 R. Michaels production started
  • 324450 06/04/10 07:59 Sereres John Shift Summary
  • 324449 06/04/10 07:57 C. Gu Target Screen
  • 324448 06/04/10 07:56 jwexler New HVs for detectors (120Hz, PbD#3)
  • 324447 06/04/10 07:46 adaq HV error message
  • 324446 06/04/10 07:42 Fansler Magnet problems last night
  • 324442 06/04/10 07:14 C. Gu when we turn on the 50uA beam at 6:10
  • 324429 06/04/10 06:04 R. Michaels getting into production, part 1
  • 324423 06/04/10 05:53 R. Michaels reference sieve pattern
  • 324422 06/04/10 05:06 R. Michaels back out of counting mode.
  • 324421 06/04/10 05:03 R. Michaels sieve slit runs
  • 324409 06/04/10 04:33 R. Barbieri gems
  • 324408 06/04/10 04:27 C. Gu Target move to BeO
  • 324407 06/04/10 04:23 R. Michaels sieves put in
  • 324406 06/04/10 04:00 R. Michaels prelim. Qsq
  • 324403 06/04/10 03:55 R. Michaels R-HRS alignment check
  • 324402 06/04/10 03:50 R.Michaels L-HRS alignment check
  • 324401 06/04/10 03:44 R. Michaels Qsq runs, part 1
  • 324400 06/04/10 03:43 A. Camsonne GEM HV changed
  • 324399 06/04/10 03:35 A. Camsonne GEM pedestal run updated
  • 324394 06/04/10 03:25 jwexler Counting mode HVs set
  • 324387 06/04/10 03:14 R. Michaels spot++ with Pb/D 3
  • 324383 06/04/10 03:06 R. Barbieri spot++, T5
  • 324379 06/04/10 03:02 A. Camsonne Left Q1 recovered
  • 324378 06/04/10 03:01 C. Gu Move Target to Pb Diamond #3
  • 324376 06/04/10 02:59 R. Barbieri spot++, BeO, T5
  • 324374 06/04/10 02:52 A. Camsonne Left arm Q1 quenched
  • 324372 06/04/10 02:39 C. Gu Target move to BeO
  • 324371 06/04/10 02:35 jwexler Raster size changed
  • 324370 06/04/10 02:33 jwexler Second harp scan complete finished, but with problems - moved on to raster check
  • 324369 06/04/10 02:31 R. Barbieri SPOT++
  • 324367 06/04/10 02:22 A. Camsonne Spot++ run 27717
  • 324364 06/04/10 01:58 jwexler LATE ENTRY: Harp scan looks off, spot small
  • 324363 06/04/10 00:43 jwexler Requested harp scan
  • 324361 06/04/10 00:20 A. Camsonne Compton dipole non turning on - Requested to go straight
  • Thursday

  • 324360 06/03/10 23:27 Asenath ETIL Hall A Tools
  • 324359 06/03/10 23:25 Kiadtisak Target Screen
  • 324344 06/03/10 22:02 A. Camsonne Septa ramped up at 729 A
  • 324343 06/03/10 21:22 A. Camsonne Q3 signs for a few different period of time
  • 324342 06/03/10 21:03 Kiadtisak Target moved to empty target
  • 324341 06/03/10 20:47 ZAFAR AHMED Left HRS run 27716 is GEM pedestal run.
  • 324339 06/03/10 20:45 A. Camsonne Polarity of right Q3
  • 324338 06/03/10 20:41 A. Camsonne Sweep in progress
  • 324334 06/03/10 20:38 A. Camsonne Q2 and Q3 cycled on both spectrometers
  • 324333 06/03/10 20:34 M Friend Compton photon detector lead shielding
  • 324332 06/03/10 20:30 hsmith Right Dipole
  • 324329 06/03/10 20:12 hsmith Left Q3
  • 324328 06/03/10 20:11 A. Camsonne ROC31 g0inj crashed again
  • 324324 06/03/10 20:04 A. Camsonne Rebooted g0inj crate
  • 324323 06/03/10 20:04 Kiadtisak GEM HV changed
  • 324315 06/03/10 19:38 Kiadtisak GEM HV changed
  • 324309 06/03/10 19:04 A. Camsonne Left Q3 back
  • 324308 06/03/10 18:36 A. Camsonne [Techs] Right dipole waiting for reset
  • 324307 06/03/10 18:01 A. Camsonne [TECH] Left HRS Q3 quench fault does not clear
  • 324306 06/03/10 17:25 jpchen Target alarm limits set
  • 324305 06/03/10 17:23 A. Camsonne Magnets in quench fault
  • 324304 06/03/10 17:20 jpchen target cool-down completed
  • 324303 06/03/10 17:19 A. Camsonne Magnets status
  • 324302 06/03/10 17:07 jpchen switched to cold return
  • 324301 06/03/10 16:48 A. Camsonne MCC update - Beam delivery should start around 8 PM
  • 324300 06/03/10 16:31 jwexler Helicity board reset to 120 Hz
  • 324299 06/03/10 16:01 jwexler ALL detectors in counting mode
  • 324297 06/03/10 15:37 jpchen Starting open J-T valve
  • 324292 06/03/10 15:13 R. Michaels present raster frequencies
  • 324286 06/03/10 14:54 M Friend GEM database updated
  • 324285 06/03/10 14:38 A. Camsonne Aluminum covers for Left Arm VDC off
  • 324281 06/03/10 14:30 R. Michaels problems with adaql2 X server (workaround in place)
  • 324280 06/03/10 14:30 jwexler Right arm in counting mode
  • 324279 06/03/10 14:29 jpchen target cooldown started.
  • 324270 06/03/10 14:25 R. Michaels prep work for data taking
  • 324266 06/03/10 13:49 A. Camsonne Target cool down procedure started
  • 324264 06/03/10 13:42 O.Glamazdin Moller magnet has been filled
  • 324261 06/03/10 13:14 Jessie B. Check list / Tech
  • 324249 06/03/10 11:49 Abdurahim Compton Cavity is fine
  • 324243 06/03/10 11:01 ZAFAR AHMED Target moved to BeO
  • 324239 06/03/10 10:34 A. Camsonne Experiment - Hurricane preparedness test update
  • 324234 06/03/10 09:51 A. Camsonne Target Vacuum status - Good for cooldown
  • Wednesday

  • 324210 06/02/10 16:58 folts target vacuum
  • 324185 06/02/10 13:51 jwexler Left Arm detectors at -2000V & scintillators ON
  • 324163 06/02/10 10:22 fansler RQ1 N2
  • 324162 06/02/10 10:02 O.Glamazdin Moller magnet has been filled
  • 324158 06/02/10 09:38 A. Camsonne Chamber being surveyed and leak checked - Cool down most likely postponed
  • Tuesday

  • 324143 06/01/10 22:13 ZAFAR AHMED Right Arm VDCs HV is Off
  • 324137 06/01/10 21:34 ZAFAR AHMED HV of VDCs is ON
  • 324130 06/01/10 19:41 jwexler Upper scintillator ON
  • 324115 06/01/10 16:47 folts target vacuum/tech
  • 324061 06/01/10 10:20 mercado Helicity @ 240Hz